### v1

# Data-Driven Goal Recognition Design for General Behavioral Agents

Abstract
Goal recognition design aims to make limited modifications to decision-making environments with the goal of making it easier to infer the goals of agents acting within those environments. Although various research efforts have been made in goal recognition design, existing approaches are computationally demanding and often assume that agents are (near-)optimal in their decision-making. To address these limitations, we introduce a data-driven approach to goal recognition design that can account for agents with general behavioral models. Following existing literature, we use worst-case distinctiveness (wcd) as a measure of the difficulty in inferring the goal of an agent in a decision-making environment. Our approach begins by training a machine learning model to predict the wcd for a given environment and the agent behavior model. We then propose a gradient-based optimization framework that accommodates various constraints to optimize decision-making environments for enhanced goal recognition. Through extensive simulations, we demonstrate that our approach outperforms existing methods in reducing wcd and enhancing runtime efficiency in conventional setups, and it also adapts to scenarios not previously covered in the literature, such as those involving flexible budget constraints, more complex environments, and suboptimal agent behavior. Moreover, we have conducted human-subject experiments which confirm that our method can create environments that facilitate efficient goal recognition from real-world human decision-makers.
1 Introduction
With the rapid advancement of artificial intelligence (AI), there has been a surge in interest in human-AI collaboration, aiming to synergize human and AI capabilities across various domains such as gaming, e-commerce, healthcare, industrial engineering, and workflow productivity. Designing AI agents to work alongside humans requires these agents to understand and infer human goals and intentions. While there has been abundant research in goal recognition (Ramírez and Geffner, 2010; Sukthankar et al., 2014) that aims to infer human goals by observing their actions, this work focuses on the goal recognition design problem (Keren et al., 2014), where one needs to identify how to modify decision-making environments to enable better goal recognition.
Our work is built on the goal recognition design problem formulated by Keren et al. (2014). In their work, they proposed the worst-case distinctiveness (wcd) metric, defined as the maximum number of decisions an agent can make without revealing its goal, to measure the difficulty of inferring the agent’s goal. They then aimed to modify the decision-making environment, through removing allowable actions from states, to optimize this measure. Since the introduction of this work, there have been several follow-up works that extend the formulation to deal with different settings, such as stochastic settings (Wayllace et al., 2016, 2017) and partial observability (Keren et al., 2016). More discussion can be found in the survey by Keren et al. (2021).
While there has been significant progress in goal recognition design, there are two main limitations in the literature. First, current approaches require evaluating the difficulty of goal recognition, i.e., worst-case distinctiveness (wcd), for a large number of potential modifications to the decision-making environment. This process is computationally demanding and limits the scalability of these approaches, especially in virtual domains where the number of possible environment configurations is large. Second, decision-making agents in goal recognition design are often assumed to be optimal. While there are efforts addressing suboptimal agents (Keren et al., 2015; Wayllace and Yeoh, 2022), these primarily focus on settings where there is a limited number of deviations from optimal decision-making. This assumption of optimality is limiting in scenarios involving human agents, who are known to often systematically deviate from optimal decision-making due to cognitive and informational constraints (Kahneman, 2003; O’Donoghue and Rabin, 1999).
To address these limitations, we propose a framework for data-driven goal recognition design with general agent behavior. We explicitly incorporate models of agent behavior into the optimization framework to better represent behavioral agents. To tackle the computational challenges, our approach leverages data-driven methods for goal recognition design. The core idea involves building a machine learning oracle that predicts the difficulty of goal recognition (e.g., wcd) given a decision-making environment and an agent’s behavioral model. This oracle is trained on datasets generated from simulations, allowing for the evaluation of wcd in any given environment and agent model. Such an approach significantly accelerates the evaluation of wcd during run-time optimization. Once the machine learning oracle is established, we apply gradient-based optimization to minimize wcd, employing Lagrangian relaxation to manage constraints on environment modifications.
To evaluate our data-driven goal recognition design framework, we conducted a series of simulations. Initially, these simulations were carried out in the standard setup in the grid world envrionment assuming optimal agent behavior. We show that our approach outperforms existing baselines in reducing worst-case distinctiveness (wcd) and demonstrates considerably better run-time efficiency. We then conducted additional simulations to showcase that our approach can generalize to settings beyond the capabilities of existing approaches in the literature. These include scenarios involving flexible budget constraints, more complex environments, and suboptimal agent behavior. Lastly, we have conducted human-subject studies demonstrating that our method can be leveraged to design environments that facilitate efficient goal recognition from real-world human decision-makers. The results highlight the potential of our approach to enable more efficient human-AI collaboration.
Contributions.
The main contributions of this work can be summarized as follows.
  • $\bullet$
    We propose a data-driven optimization framework for goal recognition design. This framework comprises a predictive model that estimates the wcd for a given environment and a model of agent behavior. Leveraging the differentiable property of the predictive model, we utilize a gradient-based optimization method for goal recognition design. The framework is flexible to accommodate various environments, design spaces, and agent models.
  • $\bullet$
    Through extensive simulations, we show that our framework not only outperforms existing approaches in goal recognition design within standard settings but also adapts to scenarios that existing approaches cannot handle. These scenarios include those with general optimization criteria, complex environments, and suboptimal agent behavior.
  • $\bullet$
    Through real-world human-subject experiments, we demonstrate that our approach can adapt to agent models trained on human behavior. Furthermore, our framework can create environments that enable more effective goal recognition with real-world human decision-makers.
2 Related Work
Our work contributes to the expanding field of human-AI collaboration. Recent research has indicated that optimizing AI alone is insufficient for maximizing the performance of human-AI teams (Bansal et al., 2021; Carroll et al., 2019). To develop truly collaborative AI agents, it is crucial to equip them with the ability to comprehend and predict the intentions and goals of their human counterparts. This challenge lies at the heart of goal recognition research (Kautz and others, 1991; Ramírez and Geffner, 2010; Hong, 2001; Sukthankar et al., 2014; Pereira et al., 2017; Masters et al., 2021). Our work specifically focuses on goal recognition design, an extension of goal recognition that includes modifying environments to better facilitate the process of recognizing goals.
Goal recognition design was originally formulated by Keren et al. (2014). Since this seminal work, numerous research efforts have expanded the concept to accommodate stochastic environments (Wayllace et al., 2017; Wayllace and Yeoh, 2022), different levels of observability (Keren et al., 2016; Wayllace et al., 2020), and a variety of design spaces (Mirsky et al., 2019). The studies most closely aligned with our approach focus on suboptimal agents (Keren et al., 2015; Wayllace and Yeoh, 2022). However, these studies characterize suboptimality by limiting deviations from the optimal policy, a method that may not adequately represent behavior of human agents, who frequently deviate systematically from optimal decision-making. Additionally, most existing work requires evaluating the difficulty of goal recognition in run time for numerous environmental modifications, which constrains the scalability of these methods. Our work sets itself apart by broadening the goal recognition design question to incorporate general agent behavioral models and by implementing a data-driven optimization approach.
From a technical standpoint, we adopt a data-driven optimization approach for goal recognition design. The use of data-driven tools in optimization has gained increasing prominence in the field of mechanism design, as evidenced by a range of recent studies (Dütting et al., 2019; Feng et al., 2018; Golowich et al., 2018; Curry et al., 2020; Kuo et al., 2020; Rahme et al., 2020; Peri et al., 2021; Cornelisse et al., 2022; Yu et al., 2023). In this regard, our work closely aligns with the studies by Yu and Ho (2022) and Yu et al. (2023), who explicitly encoded human behavior and responses within their mechanism designs for optimizing the information policy in environment and information design. While our work shares a similar motivation, our approach and problem formulation differ from theirs.
3 Problem Formulation and Methods
3.1 Problem Setting
We formulate the decision-making environment and behavioral agent models. We also explain the commonly-used metrics for evaluating goal recognition difficulty.
Decision-making environment.
We define the decision-making environment as a Markov Decision Process (MDP), represented by $W=\langle S,A,P,R\rangle$. Here, $S$ denotes the set of states, $A$ represents the set of agent actions, $P(s^{\prime}|s,a)$ is the transition probability from state $s$ to state $s^{\prime}$ upon taking action $a\in A$, and $R(s,a,s^{\prime})$ represents the bounded reward received after taking action $a$ in state $s$ and reaching state $s^{\prime}$. To emphasize the goal recognition aspects of the problem, we introduce a set of goal states $G\subseteq S$. These goal states are terminal; that is, $P(g|g,a)=1$ $\forall g\in G,a\in A$.
Models of behavioral agents.
We represent the agent’s decision-making policy in a general form $\Pi:S\times T\rightarrow A$. Specifically, for an agent with a decision-making policy $\pi\in\Pi$, the agent will execute the action $\pi(s,t)$ in state $s$ at time $t$. The agent is conceptualized as a planner $H:W\rightarrow\Pi$, where the input is an environment $w\in W$, and the output is a policy $\pi\in\Pi$. To illustrate our formulation, consider a agent parameterized by a time-variant discounting function $d(t)$. The standard optimal agent model corresponds to a fixed discounting factor $\gamma\in(0,1]$ with $d(t)=\gamma^{t}$. We can also represent an agent with present bias (O’Donoghue and Rabin, 1999) by adopting a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$, where $k>0$. We would like to highlight that our approach not only accounts for standard analytical closed-form expression of agent behavior. It can also accounts for scenarios where an agent policy $\pi$ is a machine learning model, i.e., a neural network trained on human behavioral data.
Worst-case distinctiveness (wcd).
To evaluate the difficulty of goal recognition, we adhere to standard literature and focus on the measure of worst-case distinctiveness (wcd(Keren et al., 2014), defined as the maximum number of steps an agent can take before revealing its goal. To compute the wcd for a given suboptimal agent $h\in H$ in an environment $w\in W$, we evaluate the path for the agent to each goal and compute the number of actions from the initial state that are identical for every goal.
3.2 Goal Recognition Design with General Behavioral Agents
We formalize the goal recognition design problem with general behavioral agents. Given an environment $w$ and an agent with a behavior model $h$, we denote the worst-case distinctiveness of environment $w$ for agent $h$ as $\textit{wcd}(w,h)$. Each type of modification $1\leq i\leq N$ will incur a cost $c_{i}(w,w^{\prime})$ that must fall in budget constraint $B_{i}$. The objective of the goal recognition design problem is to alter the environment from $w$ to $w^{\prime}$ in a way that minimizes $\textit{wcd}(w^{\prime},h)$, while satisfying the constraint that the cost of the modifications does not exceed the budget.
$\underset{w^{\prime}}{\text{minimize}}$ $\textit{wcd}(w^{\prime},h)$ (1)
subject to $c_{i}(w,w^{\prime})\leq B_{i},\forall 1\leq i\leq N$
3.3 Our Proposed Method
Existing approaches in goal recognition design require evaluating wcd for a large number of potential modifications to the environments to identify the optimal modifications. Since evaluating wcd requires evaluating the agent policy for multiple goals, it presents significant computational overhead and limits the scalability of these approaches. To overcome this challenge, we propose leveraging machine learning to expedite run-time computations. The main idea, as summarized in Figure 1, is to first train a machine learning model that predicts wcd for any given pair of decision-making environment $w$ and agent behavioral model $h$. After obtaining this machine learning model and utilizing its differentiable property, we develop an optimization framework that applies gradient-based optimization methods to the Lagrangian relaxation of the constrained optimization formulation in (1).
(a) Predictive model for wcd.
(b) Optimization framework.
Figure 1: We first train a predictive model to predict wcd from simulated data. We then perform gradient-based optimization that leverages the predictive model to identify environment modifications that minimize wcd with a given model of agent behavior.
Predictive model for wcd.
To build the predictive model for wcd, we first curate a training dataset through simulations. For a given environment $w$ and agent behavioral model $h$, we can obtain $\textit{wcd}(w,h)$ by solving for the agent’s actions towards different goals. After collecting a training dataset, we train the predictive model using a convolutional neural network. Implementation details are included in the appendix.
Optimization procedure.
After obtaining the predictive model, we develop a gradient-based optimization framework. Note that this optimization framework strictly generalizes the existing literature in goal recognition design where the space of modifications are limited (e.g., usually limits to blocking an action in MDP) (Keren et al., 2021).
The first step involves transforming the constrained optimization problem in (1) into an unconstrained optimization problem using Lagrangian relaxation:
$\mathcal{L}=\textit{wcd}(w^{\prime},h)+\sum_{i}\lambda_{i}(c_{i}(w,w^{\prime})-B_{i}).$
We then perform gradient descent on the relaxed Lagrangian. As environment modifications are often discrete (e.g., whether to block a cell in the grid world), we apply a discrete gradient descent procedure. Specifically, at each step of gradient descent, we obtain a gradient, which is a vector indicating the suggested change magnitude for each element (such as a cell in the grid world). We then select the element with the highest gradient value and make the corresponding change. Note that some suggested modifications may not be valid; for instance, we cannot block a cell that is already blocked in the grid world. In such cases, we proceed to the element with the next highest gradient value, continuing this process until a valid modification is made. This modification procedure is repeated until the gradient descent converges.
Note that with this Lagrangian relaxation, we cannot directly set the budget, however, based on duality, a larger Lagrangian multiplier $\lambda$ corresponds to a smaller budget $B$, in the original constrained optimization formulation. In our experiments, instead of selecting different budgets directly, we choose varying Lagrangian multipliers and record the realized costs of the modifications.
4 Experiment Setup
4.1 Benchmark Domains
We utilize two benchmark domains. The first is the standard basic grid world environment, commonly used in goal recognition design literature. The second is the Overcooked-AI environment (Carroll et al., 2019), an complex environment with a richer set of environment modification. This environment is particularly relevant to the downstream implications of our work, namely in supporting human-AI collaboration.
(a) Grid world.
(b) Overcooked-AI.
Figure 2: The example showcases two benchmark environments. The first, displayed on the left, is a grid world environment. The agent starts at a position marked ’S’ and aims to reach one of the goal positions labeled ’G’. The agent must navigate through the grid, avoiding blocked cells marked with ’x’. The second environment, illustrated on the right, is an Overcooked-AI setting. In this scenario, the agent’s objective is to pick up ingredients and complete the cooking of their target recipe, which constitutes their goal.
4.1.1 Benchmark Domain: Grid world
In the grid world domain, agents navigate a grid with several potential goals. Take, for example, a grid world environment as illustrated in Figure 2(a). In this environment, an optimal agent starts at point ’S’ and aims for one of the goals. Spaces marked ’X’ are blocked, barring the agent’s passage. In this particular environment, the worst-case distinctiveness (wcd) is 0, as an optimal agent, following the shortest path to either goal, would reveal its intended goal on the first move due to non-overlapping paths. Our experiments are primarily focused on grid world environments with two goals for simplicity. However, our approach is applicable to scenarios with more than two goals.
Design space of environment modifications.
In the context of goal recognition design in grid world, the space environment modification is often limited to adding blocks to spaces (Keren et al., 2014), also called action removal in the literature. In our work, we broaden the design space for modifications to also consider the unblocking of existing blocked spaces as potential modifications.
4.1.2 Benchmark Domain: Overcooked-AI
We also conduct our evaluations on a more complex domain: Overcooked-AI [^1] https://github.com/HumanCompatibleAI/overcooked_ai. This environment is based on the popular multiplayer game Overcooked. In each game, players (agents) collaborate to prepare and deliver specific recipes. Since the goal of our approach is to enable efficient inference of agent goals, we focus on the special case with a single agent. The environment is represented as a grid (see Figure 2(b)) with each cell specifying the object that is placed at the cell. The objects may be a counter, a tomato, an onion, a pot, a dish, a serving point, or an empty space. The agent can only occupy empty spaces in the environment and cannot step on any other objects in the environment. The agent can carry the movable objects, i.e. onion, tomato, and dish, and drop them on any of the other non-space objects (counter, serving point, or pot). To navigate in the environment, the agent can move left, right, north, or south and maintain an orientation that is consistent with the last movement direction.
Goal recognition in Overcooked-AI.
The goal of the agent in Overcooked-AI is defined by the recipe it needs to complete. For instance, to prepare a soup with one onion and two tomatoes, the agent must pick up the ingredients, place them in the pot, and cook them. In the context of goal recognition, the objective is to deduce which recipe the agent is preparing. For the sake of simplicity in our experiments, we focus on scenarios with only two possible goals or recipes.
Design space of environment modifications.
A primary challenge in the Overcooked-AI environment is its considerably larger design space for environment modifications. The design space includes the addition, removal, or changing position of any object within the environment, subject to the modification being valid. This means an object cannot be added to a cell already occupied by another object, and an object that does not exist cannot be removed. The objective of conducting experiments in this domain is to assess whether our approach can be used to address more complex domains.
4.2 Baselines
We now describe the baseline approaches we use to compare against our approach.
  • $\bullet$
    Exhaustive search: This is the brute-force approach that evaluates wcd for all the environments on the search path until the minimum possible wcd is found. It is guaranteed to achieve minimum wcd. However, given the computation overhead, this approach is not applicable in most situations.
  • $\bullet$
    Pruned-Reduce (Keren et al., 2014): This baseline is specifically designed for settings where modifications are limited to action removal such as blocking a cell in grid world. It speeds up the exhaustive search and retains the optimal property. However, its scalability is still limited.
  • $\bullet$
    Greedy search using true wcd: This greedy search baseline finds the single environment modification that leads to the maximum reduction of wcd at each iteration. This approach requires to evaluate the wcd for all possible single environment modifications at each iteration.
  • $\bullet$
    Greedy search using predicted wcd: In addition to greedy search using true wcd, we leverage our predictive model for wcd and design another greedy baseline. This baseline finds the single environment modification that leads to the maximum reduction of predicted wcd at each iteration.
4.3 Models of Agent Behavior
In our experiments, to demonstrate that our approach works for different models of agent behavior, we have examined three types of agent behavior.
  • $\bullet$
    Optimal agent behavior. The first one is the standard optimal agent behavior. Conducting experiments with optimal agent behavior enables us to compare our approach with standard approaches in the literature, which are often developed under the optimal agent assumption.
  • $\bullet$
    Parameterized suboptimal agent behavior. We also consider an generalized behavior model that is captured by a time-variant discounting function $d(t)$, and the agent’s objective is to optimize a time discounted reward with the discounting factor for $t$ steps in the future being $d(t)$. The standard optimal agent model corresponds to a fixed discounting factor $\gamma\in(0,1]$ with $d(t)=\gamma^{t}$. We can also represent an agent with present bias (O’Donoghue and Rabin, 1999) by adopting a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$, where $k>0$.
  • $\bullet$
    Data-driven agent behavior. We also address settings where the model of agent behavior is a machine learning model trained on human behavioral data.
5 Experiments
We have conducted extensive simulations to examine the performance of our approach. Moreover, we have conducted real-world human-subject experiments to assess whether our approach is effective with human decision-makers.
5.1 Simulations
In our simulations, we first examine how our approach compares with existing methods in standard setups found in the literature. We then showcase the generalizability of our approach by extending it to scenarios that go beyond the setups studied in the literature, including scenarios with more dynamic budget constraints, more complicated environments, and suboptimal agent behavior.
5.1.1 Standard setting
We first focus on settings within the grid world domain, where modifications are limited to blocking cells and agent behavior is assumed to be optimal. This scenario forms the standard setting for the majority of goal recognition design studies, as highlighted in Table 1 of the survey by Keren et al. (2021). The objective of this set of simulations is to enable comparisons with state-of-the-art methods.
In our experiments, the initial environment is generated randomly: we first randomly select the number of blocked cells from the range $[0,12]$, followed by the random allocation of the blocked cells’ positions. We also randomly determine the starting grid and two goal grids. Environments where the goals are not reachable from the starting grid are filtered out. Following this procedure, we randomly generate 200 environments and compare the average performance of our approach with that of baseline methods.
(a) $6\times 6$ grids
(b) $13\times 13$ grids
Figure 3: wcd reduction in a grid world when only blocking modifications are allowed. Exhaustive search and Pruned-Reduce are not included in the larger environments of (b) because they cannot complete the computation within an hour for a single environment.
We start our experiments with a grid world of size $6\times 6$. In this simplest scenario, our approach and all baselines exhibit similar performance in wcd reduction, as depicted in Figure 3(a). For runtime comparison, our method demonstrates a significant speed advantage over exhaustive search, performing on average 10 times faster. With the maximum allowed budget of 19 modifications, our approach required only 0.2 seconds, whereas the exhaustive search took approximately 2 seconds. [^2] Detailed runtime comparisons are included in the appendix. In this simplest setting, the runtime of our approach roughly matches that of the greedy method but is 10 times faster than exhaustive search. In all other settings, our method is several orders of magnitude faster than both exhaustive search and greedy using true wcd. Although the greedy baseline becomes much faster when utilizing our predicted model for wcd, our approach remains over 3 times faster than the greedy baseline with predicted wcd for large budgets. We then expanded the grid size to $13\times 13$. In this more complex setting, both exhaustive search and the Pruned-Reduce method were unable to complete computations within an hour for an instance, leading to their exclusion from the baselines. As illustrated in Figure 3(b), our approach surpasses the greedy baselines in wcd reduction and is 3 times faster than these baselines for large budgets.
5.1.2 Settings with flexible budget constraints
In the literature, most works focus on a single type of environment modification (e.g., blocking a cell). Given the flexibility of our optimization framework, we extend our approach to also include ’unblocking’ blocked cells as a possible environment update. In our simulations, we examine two common cases. In the shared budget setting, the total number of blocking and unblocking actions is bounded by a given shared budget. In the individual budget setting, the number of blocking and unblocking actions is bounded by different budgets. Specifically, we allow the number of blocking actions to be 5 times the number of unblocking actions (rounded to the nearest integer). Given that there are no established baselines for this setting in the literature, we compare our results against greedy baselines. The results, as shown in Figure 4, demonstrate the effectiveness of our approach in wcd reduction. Regarding runtime, our approach is several orders of magnitude faster than the greedy method with true wcd and at least 3 times faster than the greedy method with predicted wcd.
(a) Shared budget.
(b) Individual budget.
Figure 4: wcd reduction in settings with two types of modifications.
5.1.3 Complex domain and suboptimal agent behavior
We consider two additional extensions to the standard setting. In the first, we evaluate the performance of our approach in a more complex problem domain: Overcooked-AI. In the second, we return to the grid world but include scenarios with suboptimal agent behavior. Both extensions utilize a grid size of $6\times 6$. For the Overcooked-AI environment, we assume optimal agent behavior and aim to explore how our approach adapts to a much richer space of environment modifications. For the suboptimal human behavior, we employ the model described in Section 4.3, utilizing a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$ and set $k=8$. For both extensions, standard approaches such as exhaustive search and Prune-Reduce are either too slow or not applicable. Therefore, we compare our results with the greedy baselines. The results for both extensions, as illustrated in Figures 6 and 6, demonstrate that our approach adapts well to both settings. Regarding runtime, our approach is again several orders of magnitude faster than the greedy method with true wcd and at least 3 times faster than the greedy method using predicted wcd.
Figure 5: Overcooked-AI Environment. Assuming optimal agent behavior.
Figure 6: Grid world of size $6\times 6$. Incorporating suboptimal agent behavior.
5.2 Real-World Human-Subject Experiments
In our simulations, we demonstrate that our approach consistently outperforms baseline methods in terms of wcd reduction and offers greater efficiency in runtime. To assess the applicability of our approach in settings where humans are the decision-makers, we conducted two sets of experiments involving human subjects. In the first experiment, our goal is to collect human behavioral data within our decision-making environments. Utilizing this data, we employ imitation learning to develop a model that accurately represents human behavior. This model is then integrated as the agent behavior model within our approach. In our second experiment, we aim to evaluate whether our approach indeed leads to environments that facilitate more effective goal recognition by human decision-makers. These experiments are approved by the Institutional Review Board (IRB) of our institution.
5.2.1 Experiment 1: Collection of human behavioral data
In our experiment, we recruited participants to play 15 navigation games within a $6\times 6$ grid world. In each game, participants were tasked with navigating from a start position to a designated goal. At each time step, they could choose to move in one of four directions: Up, Down, Right, or Left. A game concluded when the participant reached the goal. The environments for these games were generated similarly to our simulations, with start positions, goal positions, and block positions all being randomly determined. Our objective with this setup was to leverage the collected data to develop a data-driven model of human behavior. We recruited 200 workers from Amazon Mechanical Turk for this study, paying an average hourly rate of approximately \$16 per worker.
Figure 7: The interface of our human subject experiment.
Table 1: The prediction accuracy for human behavior assuming optimal behavior and using data-driven model in Experiment 1.
Training Accuracy Validation Accuracy Testing Accuracy
Assuming Optimal Behavior 0.7266 0.6964 0.7131
Data-Driven Model 0.9189 0.8136 0.8422
Learning models of human behavior.
The collected human data were divided into training, validation, and testing sets: the training set included data from 160 workers, comprising approximately 70,000 instances of user decisions, while the validation and testing sets each contained data from 20 workers, amounting to around 8,800 instances of user decisions each. We employed a 4-layer Multilayer Perceptron (MLP) model for training the model. The input to the model is the current environment layout, and its output predicts the next human action. We fine-tuned the hyperparameters, including learning rate, hidden layer size, and L2 penalty, based on the performance on the validation dataset.
We compared the performance of our learned model against a model that assumes optimal agent behavior, which is defined as taking the shortest path to the goal. The training, validation, and test accuracies of both models are presented in Table 1. These results clearly indicate that human behavior deviates significantly from optimality. This deviation underscores the importance of incorporating a realistic model of human behavior in goal recognition design, particularly when humans are the decision-makers.
5.2.2 Experiment 2: Evaluating goal recognition design
We next evaluate the performance of our approach that incorporates the data-driven model of human behavior from experiment 1. To do this, we randomly generate 30 initial environments using the same setup as in our simulations. These environments are then updated according to four different methods, all operating within a modification budget of 20.
  • $\bullet$
    Initial: No updates to the environment.
  • $\bullet$
    Greedy: Greedy baseline using predicted wcd from the data-driven human behavior model.
  • $\bullet$
    Proposed (opt-bhvr): Our proposed approach when assuming the agent is following optimal behavior (i.e., picking one of the shortest paths towards the goal).
  • $\bullet$
    Proposed (data-driven): Our proposed approach using predicted wcd from the data-driven model.
We recruited 200 workers from Amazon Mechanical Turk. Each worker was randomly assigned to one of the four treatments specified above, with the distinction between treatments being the environments presented to them. Workers were randomly assigned a goal for each environment and tasked with navigating to reach it. After collecting their series of actions aimed at reaching the goal, we analyzed this behavioral data to assess the effectiveness of goal recognition design for each approach.
Comparing average overlapping path.
To evaluate the effectiveness of each goal recognition design approach, we first measure the average overlapping path towards each of the two goals. Specifically, each recruited worker is exposed to all 30 environments in their assigned treatment and is instructed to reach one of two randomly selected goals. For each treatment, we calculate the number of overlapping actions for every pair of workers assigned to different goals within each environment. We then compute the average of these numbers, yielding the average overlapping path as our performance measure. This metric reflects the difficulty of inferring agent goals for the envioronment and serves as an empirical proxy for wcd. With this metric, we vary the initial $k$ portion of actions taken by the workers in our analysis and observe how the average overlapping path changes with the amount of observed actions, denoted by $k$. As depicted in Figure 10(c), our approach, when integrated with data-driven models, leads to environments that facilitate the easier identification of the humans’ intended goal.
Comparing the accuracy of goal inference.
We next measure the accuracy of inferring human goals based on partial observations, utilizing an off-the-shelf Bayesian inference algorithm [^3] We assume an initial uniform prior distribution. As the agent takes actions, we update the posterior belief over goals based on the likelihood probability of the data-driven behavior model prediction. . Specifically, for each worker, we reveal the first $k$ portion of the worker’s actions to the inference algorithm to infer the worker’s goal. This enables us to compute the average inference accuracy. The results, as shown in Figure 8(b), demonstrate that our approach leads to environments that are easier for goal recognition.
(a) Averaging overlapping path (the lower the better.)
(b) Bayesian goal inference (the higher the better).
Figure 8: Comparing different goal recognition design approaches in our human subject experiments. Our approach coupled with data-driven models is shown to generate environments that enable the most effective goal recognition with real human decision makers.
6 Conclusion
Effective collaboration between humans and AI hinges on the AI’s ability to infer intentions and goals of humans. In this study, we work on the problem of goal recognition design, aiming to update the decision-making environments to make it easier for the AI agents to perform goal recognition. By addressing the key limitations in the existing literature, notably the computational demand and the assumption of optimal agent behavior, we have developed a data-driven optimization framework that is both efficient and scalable. Moreover, our approach is capable of integrating general models of various agent behavior, moving beyond the traditional assumption of optimality and making it more applicable to real-world scenarios involving human agents.
Future work of our study includes extending our framework to more complex environments, particularly those that are dynamic and partially observable. Given our framework is adaptable to different forms of agent behavior, understanding how different patterns of agent suboptimality impacts the difficulty of goal recognition could lead to useful insights for enabling better human-AI collaboration.
Appendix A Experiment Details and Additional Results
In this section, we provide details of the experiments that have not been included in the main paper due to space constraints. We also include and discuss additional experiment results.
A.1 Implementation Details
In this subsection, we discuss the details of our implementation for training the predictive model for wcd. To train a CNN-based prediction model, we generated an extensive training dataset by solving wcd for a large number of randomly selected environments, given a specific agent model. For creating a predictive model for environments with a grid size of $N$ by $N$, we utilized a custom Convolutional Neural Network (CNN) architecture. This architecture is tailored to process input data, $w_{i}$, consisting of $k$ $N\times N$ channels, where $k$ represents the number of potential objects in the environment.
The architecture comprises two distinct blocks: the initial block contains three successive convolutional layers, each followed by batch normalization and a rectified linear unit (ReLU) activation function. The subsequent block introduces pooling operations and additional convolutional layers for more sophisticated feature extraction. After the convolutional blocks, the architecture includes fully connected layers, leading to a single output unit. The design integrates Leaky ReLU, ReLU, and dropout layers to add non-linearity and provide regularization. This flexible architecture can adapt to tasks with multi-channel input data, allowing for modifications based on the specific needs of the task and dataset characteristics. Batch normalization and dropout are incorporated to promote training stability and prevent overfitting. We employed the Adam optimizer and Mean Squared Error (MSE) loss for training the models, specifically focusing on settings involving optimal human agents.
To create the dataset to train the model, we randomly generated $150$K grid designs and computed the corresponding wcd. This dataset was then split into the training dataset ($80\%$) and validation dataset ($20\%$). For our experiments reported in Section 5 with $13\times 13$ grid size and optimal agent behavior, we have achieved an MSE value around 0.18 for both training and validation errors. The results are shown in Figure 9. We have also examined a variety of setups, and generally we can reach small errors for the predictive model that can enable the optimization.
Figure 9: Training loss and validation errors for the predictive model for wcd with $13\times 13$ grid world and optimal human behavior.
A.2 Results for Run Time Comparisons
Due to space constraints and also that the results align with one would expect, we do not include the run time details for different approaches in the main text. We include the results here for completeness. Overall, as shown in Figure 10, approaches that leverage the predictive model for wcd are orders of magnitude faster than methods that require to evaluate wcd during run-time. Note that the y-axis of the figure is in log scale so the difference is in at least two orders of magnitude.
(a) Standard Setup (small)
(b) Standard Setup (large)
(c) Suboptimal
(d) Overcooked-AI
Figure 10: Run time taken by each method in the different experimental conditions. In the standard setup with a small grid size, all methods except exhaustive complete within less than half of a second but only our approach scales with a larger grid size and other more complex configurations.
Appendix B Additional Experiment Results
In this section, we report the additional experiment results that are not included in the main text due to space constraints. In Section 5.1.2 of the main paper, we provided details of our performance with a large grid size. In a smaller grid world, our approach achieves comparable performance to the baselines but it significantly outperforms them in larger grid sizes as shown in Figure 11.
(a) Individual budget
(b) Shared budget
Figure 11: wcd reduction in smaller ($6\times 6$) gridworld with optimal agent behavior.
Appendix C Simulation Details
To generate our training and evaluation datasets, we randomly generated environments and kept environments that are valid, e.g., the goals are reachable, and that the objects don’t overlap. Below we provide more details for environment generation for specified grid sizes for both overcooked and grid-world domains.
C.0.1 Gridworld
In a grid world, a valid environment includes a starting position, blocked cells, and two goal positions. The starting position is randomly placed in the first column, and the goal positions are randomly placed in the last two columns. The number of blocked positions in each grid is randomly selected from a range of 00 to 2×2$\times$2 × the grid width. We discard any assignments that make the goals unreachable. For experiments with suboptimal behavior, we also randomly assigned small subgoal rewards to unblocked cells that the agent would collect on its way to the goal state. The two goal states assigned a large reward that is 10 times the largest subgoal reward.
C.0.2 Overcooked-AI
In Overcooked-AI, a valid environment includes one pot, one tomato source, one onion source, one dish source, one serving point, no open spaces at the border, and any number of open spaces and blocked cells. All objects must be reachable from the agent’s randomly assigned starting position, with the number and positions of blocked cells assigned randomly. The agent is randomly assigned any two possible goals: three tomato soups, three onion soups, or a mixed soup. Each goal has the same randomly assigned reward value. Suboptimal behavior is modeled by assigning small random subgoal rewards when adding ingredients to the pot.
Appendix D Details of Human-Subject Experiment
Lastly, we include more information about our human-subject experiments. In the human-subject experiment, each worker is asked to play a navigation game in $6\times 6$ grid world environments. The task interface is shown in Figure 12.
Figure 12: Human-Subject experiment interface.
We have recruited 200 workers from Amazon Mechanical Turk in total, and Table 2 contains the demographic information of the workers.
Table 2: Demographic information of the participants in our experiment.
Group Category Number
Age 20 to 29 84
30 to 39 76
40 to 49 26
50 or older 14
Gender Female 89
Male 110
Other 1
Race / Ethnicity Caucasian 175
Black or African-American 8
American Indian/Alaskan Native 3
Asian or Asian-American 8
Spanish/Hispanic 1
Other 5
Education High school degree 5
Some college credit, no degree 5
Associate’s degree 4
Bachelor’s degree 135
Graduate’s degree 49
Other 2

### v2

@@
 We then propose a gradient-based optimization framework that accommodates various constraints to optimize decision-making environments for enhanced goal recognition.
-Through extensive simulations, we demonstrate that our approach outperforms existing methods in reducing wcd and enhancing runtime efficiency in conventional setups, and it also adapts to scenarios not previously covered in the literature, such as those involving flexible budget constraints, more complex environments, and suboptimal agent behavior.
-Moreover, we have conducted human-subject experiments which confirm that our method can create environments that facilitate efficient goal recognition from real-world human decision-makers.
+Through extensive simulations, we demonstrate that our approach outperforms existing methods in reducing wcd and enhancing runtime efficiency in conventional setup.
+Morevoer, our approach also adapts to settings in which existing approaches do not apply, such as those involving flexible budget constraints, more complex environments, and suboptimal agent behavior.
+Finally, we have conducted human-subject experiments which confirm that our method can create environments that facilitate efficient goal recognition from real-world human decision-makers.
 1 Introduction
-With the rapid advancement of artificial intelligence (AI), there has been a surge in interest in human-AI collaboration, aiming to synergize human and AI capabilities across various domains such as gaming, e-commerce, healthcare, industrial engineering, and workflow productivity.
+With the rapid advancement of artificial intelligence (AI), there has been a surge in interest in human-AI collaboration, aiming to synergize human and AI capabilities across various domains such as gaming, e-commerce, healthcare, and workflow productivity.
 Designing AI agents to work alongside humans requires these agents to understand and infer human goals and intentions.
-While there has been abundant research in goal recognition (Ramírez and Geffner, 2010; Sukthankar et al., 2014) that aims to infer human goals by observing their actions, this work focuses on the goal recognition design problem (Keren et al., 2014), where one needs to identify how to modify decision-making environments to enable better goal recognition.
+While there has been abundant research in goal recognition [23, 24] that aims to infer human goals by observing their actions, this work focuses on the goal recognition design problem [12], where one needs to identify how to modify decision-making environments to enable better goal recognition.
 Our work is built on the goal recognition design problem formulated by Keren et al.
-(2014).
-In their work, they proposed the worst-case distinctiveness (wcd) metric, defined as the maximum number of decisions an agent can make without revealing its goal, to measure the difficulty of inferring the agent’s goal.
+[12].
+They proposed the worst-case distinctiveness (wcd) metric, defined as the maximum number of decisions an agent can make without revealing its goal, to measure the difficulty of inferring the agent’s goal.
 They then aimed to modify the decision-making environment, through removing allowable actions from states, to optimize this measure.
-Since the introduction of this work, there have been several follow-up works that extend the formulation to deal with different settings, such as stochastic settings (Wayllace et al., 2016, 2017) and partial observability (Keren et al., 2016).
+Since the introduction of this work, there have been several follow-up works that extend the formulation to deal with different settings, such as stochastic settings [28, 29] and partial observability [14].
 More discussion can be found in the survey by Keren et al.
-(2021).
+[15].
 While there has been significant progress in goal recognition design, there are two main limitations in the literature.
 First, current approaches require evaluating the difficulty of goal recognition, i.e., worst-case distinctiveness (wcd), for a large number of potential modifications to the decision-making environment.
-This process is computationally demanding and limits the scalability of these approaches, especially in virtual domains where the number of possible environment configurations is large.
+Since each evaluation of wcd requires solving the optimal policy multiple times to each of the goals, this process is computationally demanding and limits the scalability of these approaches, especially in virtual domains where the number of possible environment configurations is large.
 Second, decision-making agents in goal recognition design are often assumed to be optimal.
-While there are efforts addressing suboptimal agents (Keren et al., 2015; Wayllace and Yeoh, 2022), these primarily focus on settings where there is a limited number of deviations from optimal decision-making.
-This assumption of optimality is limiting in scenarios involving human agents, who are known to often systematically deviate from optimal decision-making due to cognitive and informational constraints (Kahneman, 2003; O’Donoghue and Rabin, 1999).
+While there are efforts addressing suboptimal agents [13, 27], these primarily focus on settings where there is a limited number of deviations from optimal decision-making.
+This assumption of optimality is limiting in scenarios involving human agents, who are known to often systematically deviate from optimal decision-making due to cognitive and informational constraints [10, 19].
 To address these limitations, we propose a framework for data-driven goal recognition design with general agent behavior.
-We explicitly incorporate models of agent behavior into the optimization framework to better represent behavioral agents.
+To relax the optimality assumption, we explicitly incorporate models of agent behavior into the optimization framework to better represent behavioral agents.
 To tackle the computational challenges, our approach leverages data-driven methods for goal recognition design.
@@
 Such an approach significantly accelerates the evaluation of wcd during run-time optimization.
-Once the machine learning oracle is established, we apply gradient-based optimization to minimize wcd, employing Lagrangian relaxation to manage constraints on environment modifications.
-To evaluate our data-driven goal recognition design framework, we conducted a series of simulations.
-Initially, these simulations were carried out in the standard setup in the grid world envrionment assuming optimal agent behavior.
+Once the machine learning oracle is established, we apply the general gradient-based optimization approach to minimize wcd, employing Lagrangian relaxation to manage constraints on environment modifications.
+This allows us to address more general forms of objectives and constraints that existing approaches in the literature cannot address.
+To evaluate our framework, we conducted a series of simulations and human-subject experiments.
+We start with simulations in the standard setup, with the grid world environment and optimal agent assumption.
 We show that our approach outperforms existing baselines in reducing worst-case distinctiveness (wcd) and demonstrates considerably better run-time efficiency.
@@
 The results highlight the potential of our approach to enable more efficient human-AI collaboration.
-Contributions.
-The main contributions of this work can be summarized as follows.
+Contributions. The main contributions of this work can be summarized as follows.
   • $\bullet$
@@
 This framework comprises a predictive model that estimates the wcd for a given environment and a model of agent behavior.
-Leveraging the differentiable property of the predictive model, we utilize a gradient-based optimization method for goal recognition design.
-The framework is flexible to accommodate various environments, design spaces, and agent models.
+We then utilize a gradient-based optimization method for goal recognition design.
+The framework is flexible to accommodate various environments, design spaces, and agent behavior.
   • $\bullet$
-    Through extensive simulations, we show that our framework not only outperforms existing approaches in goal recognition design within standard settings but also adapts to scenarios that existing approaches cannot handle.
-These scenarios include those with general optimization criteria, complex environments, and suboptimal agent behavior.
+    Through extensive simulations, we show that our framework not only outperforms existing approaches in goal recognition design within standard settings but also adapts to scenarios that existing approaches cannot handle, including general optimization criteria, complex environments, and suboptimal agent behavior.
   • $\bullet$
-    Through real-world human-subject experiments, we demonstrate that our approach can adapt to agent models trained on human behavior.
-Furthermore, our framework can create environments that enable more effective goal recognition with real-world human decision-makers.
+    Through human-subject experiments, we demonstrate that our approach can adapt to agent models trained on human behavior.
+Furthermore, our framework can create environments that enable more effective goal recognition with real-world humans.
+To the best of our knowledge, our work is the first to utilize human-subject experiments to evaluate the environment design for goal recognition.
 2 Related Work
 Our work contributes to the expanding field of human-AI collaboration.
-Recent research has indicated that optimizing AI alone is insufficient for maximizing the performance of human-AI teams (Bansal et al., 2021; Carroll et al., 2019).
+Recent research has indicated that optimizing AI alone is insufficient for maximizing the performance of human-AI teams [1, 2].
 To develop truly collaborative AI agents, it is crucial to equip them with the ability to comprehend and predict the intentions and goals of their human counterparts.
-This challenge lies at the heart of goal recognition research (Kautz and others, 1991; Ramírez and Geffner, 2010; Hong, 2001; Sukthankar et al., 2014; Pereira et al., 2017; Masters et al., 2021).
+This challenge lies at the heart of goal recognition research [11, 23, 9, 24, 20, 17].
 Our work specifically focuses on goal recognition design, an extension of goal recognition that includes modifying environments to better facilitate the process of recognizing goals.
-Goal recognition design was originally formulated by Keren et al.
-(2014).
-Since this seminal work, numerous research efforts have expanded the concept to accommodate stochastic environments (Wayllace et al., 2017; Wayllace and Yeoh, 2022), different levels of observability (Keren et al., 2016; Wayllace et al., 2020), and a variety of design spaces (Mirsky et al., 2019).
-The studies most closely aligned with our approach focus on suboptimal agents (Keren et al., 2015; Wayllace and Yeoh, 2022).
+Goal recognition design was formulated by Keren et al.
+[12].
+Since this seminal work, numerous research efforts have expanded the concept to accommodate stochastic environments [29, 27], different levels of observability [14, 30], and a variety of design spaces [18].
+The studies most closely aligned with our approach focus on suboptimal agents [13, 27].
 However, these studies characterize suboptimality by limiting deviations from the optimal policy, a method that may not adequately represent behavior of human agents, who frequently deviate systematically from optimal decision-making.
@@
 From a technical standpoint, we adopt a data-driven optimization approach for goal recognition design.
-The use of data-driven tools in optimization has gained increasing prominence in the field of mechanism design, as evidenced by a range of recent studies (Dütting et al., 2019; Feng et al., 2018; Golowich et al., 2018; Curry et al., 2020; Kuo et al., 2020; Rahme et al., 2020; Peri et al., 2021; Cornelisse et al., 2022; Yu et al., 2023).
-In this regard, our work closely aligns with the studies by Yu and Ho (2022) and Yu et al.
-(2023), who explicitly encoded human behavior and responses within their mechanism designs for optimizing the information policy in environment and information design.
+The use of data-driven tools in optimization has gained increasing prominence in the field of mechanism design, as evidenced by a range of recent studies [5, 7, 8, 4, 16, 22, 21, 3, 32].
+Moreover, our work closely aligns with the recent studies that explicitly encoded human behavior and responses in the design of computational and learning systems [25, 26, 31, 32, 6].
 While our work shares a similar motivation, our approach and problem formulation differ from theirs.
@@
 The standard optimal agent model corresponds to a fixed discounting factor $\gamma\in(0,1]$ with $d(t)=\gamma^{t}$.
-We can also represent an agent with present bias (O’Donoghue and Rabin, 1999) by adopting a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$, where $k>0$.
+We can also represent an agent with present bias [19] by adopting a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$, where $k>0$.
 We would like to highlight that our approach not only accounts for standard analytical closed-form expression of agent behavior.
@@
 Worst-case distinctiveness (wcd).
-To evaluate the difficulty of goal recognition, we adhere to standard literature and focus on the measure of worst-case distinctiveness (wcd(Keren et al., 2014), defined as the maximum number of steps an agent can take before revealing its goal.
+To evaluate the difficulty of goal recognition, we adhere to standard literature and focus on the measure of worst-case distinctiveness (wcd[12], defined as the maximum number of steps an agent can take before revealing its goal.
 To compute the wcd for a given suboptimal agent $h\in H$ in an environment $w\in W$, we evaluate the path for the agent to each goal and compute the number of actions from the initial state that are identical for every goal.
@@
 After obtaining this machine learning model and utilizing its differentiable property, we develop an optimization framework that applies gradient-based optimization methods to the Lagrangian relaxation of the constrained optimization formulation in (1).
+The main benefits of our approach compared with existing methods include: 1) the ability to incorporate different agent models $h$, 2) flexibility for various forms of optimization objectives and constraints, and 3) run-time efficiency.
 (a) Predictive model for wcd.
@@
 We then perform gradient-based optimization that leverages the predictive model to identify environment modifications that minimize wcd with a given model of agent behavior.
+(a) Grid world.
+(b) Overcooked-AI.
+Figure 2: The example showcases two benchmark environments.
+The first, displayed on the left, is a grid world environment.
+The agent starts at a position marked ’S’ and aims to reach one of the goal positions labeled ’G’.
+The agent must navigate through the grid, avoiding blocked cells marked with ’x’.
+The second environment, illustrated on the right, is an Overcooked-AI setting.
+In this scenario, the agent’s objective is to pick up ingredients and complete the cooking of their target recipe, which constitutes their goal.
 Predictive model for wcd.
-To build the predictive model for wcd, we first curate a training dataset through simulations.
+To build the predictive model for wcd, we curate a training dataset through simulations.
 For a given environment $w$ and agent behavioral model $h$, we can obtain $\textit{wcd}(w,h)$ by solving for the agent’s actions towards different goals.
@@
 After obtaining the predictive model, we develop a gradient-based optimization framework.
-Note that this optimization framework strictly generalizes the existing literature in goal recognition design where the space of modifications are limited (e.g., usually limits to blocking an action in MDP) (Keren et al., 2021).
+This framework strictly generalizes the existing literature in goal recognition design where the space of modifications are limited (e.g., usually limits to blocking an action in MDP) [15].
 The first step involves transforming the constrained optimization problem in (1) into an unconstrained optimization problem using Lagrangian relaxation:
@@
 The first is the standard basic grid world environment, commonly used in goal recognition design literature.
-The second is the Overcooked-AI environment (Carroll et al., 2019), an complex environment with a richer set of environment modification.
+The second is the Overcooked-AI environment [2], an complex environment with a richer set of environment modification.
 This environment is particularly relevant to the downstream implications of our work, namely in supporting human-AI collaboration.
-(a) Grid world.
-(b) Overcooked-AI.
-Figure 2: The example showcases two benchmark environments.
-The first, displayed on the left, is a grid world environment.
-The agent starts at a position marked ’S’ and aims to reach one of the goal positions labeled ’G’.
-The agent must navigate through the grid, avoiding blocked cells marked with ’x’.
-The second environment, illustrated on the right, is an Overcooked-AI setting.
-In this scenario, the agent’s objective is to pick up ingredients and complete the cooking of their target recipe, which constitutes their goal.
 4.1.1 Benchmark Domain: Grid world
@@
 Design space of environment modifications.
-In the context of goal recognition design in grid world, the space environment modification is often limited to adding blocks to spaces (Keren et al., 2014), also called action removal in the literature.
+In the context of goal recognition design in grid world, the space environment modification is often limited to adding blocks to spaces [12], also called action removal in the literature.
 In our work, we broaden the design space for modifications to also consider the unblocking of existing blocked spaces as potential modifications.
@@
 We also conduct our evaluations on a more complex domain: Overcooked-AI [^1] https://github.com/HumanCompatibleAI/overcooked_ai.
-This environment is based on the popular multiplayer game Overcooked.
+This environment is based on the popular game Overcooked.
 In each game, players (agents) collaborate to prepare and deliver specific recipes.
@@
 A primary challenge in the Overcooked-AI environment is its considerably larger design space for environment modifications.
-The design space includes the addition, removal, or changing position of any object within the environment, subject to the modification being valid.
-This means an object cannot be added to a cell already occupied by another object, and an object that does not exist cannot be removed.
+The design space includes the changing position of any object within the environment, subject to the modification being valid.
+This means that the change cannot result in a new design where any of the objects is unreachable by the agent.
 The objective of conducting experiments in this domain is to assess whether our approach can be used to address more complex domains.
@@
   • $\bullet$
-    Pruned-Reduce (Keren et al., 2014): This baseline is specifically designed for settings where modifications are limited to action removal such as blocking a cell in grid world.
+    Pruned-Reduce [12]: This baseline is specifically designed for settings where modifications are limited to action removal such as blocking a cell in grid world.
 It speeds up the exhaustive search and retains the optimal property.
@@
     Parameterized suboptimal agent behavior.
-We also consider an generalized behavior model that is captured by a time-variant discounting function $d(t)$, and the agent’s objective is to optimize a time discounted reward with the discounting factor for $t$ steps in the future being $d(t)$.
+We also consider an generalized behavior model parameterized by $d(t)$.
+The agent’s objective is to optimize a time discounted reward with the discounting factor for $t$ steps in the future being $d(t)$.
 The standard optimal agent model corresponds to a fixed discounting factor $\gamma\in(0,1]$ with $d(t)=\gamma^{t}$.
-We can also represent an agent with present bias (O’Donoghue and Rabin, 1999) by adopting a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$, where $k>0$.
+We can also represent an agent with present bias [19] by adopting a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$, where $k>0$.
   • $\bullet$
@@
 5 Experiments
-We have conducted extensive simulations to examine the performance of our approach.
-Moreover, we have conducted real-world human-subject experiments to assess whether our approach is effective with human decision-makers.
 5.1 Simulations
@@
 We first focus on settings within the grid world domain, where modifications are limited to blocking cells and agent behavior is assumed to be optimal.
-This scenario forms the standard setting for the majority of goal recognition design studies, as highlighted in Table 1 of the survey by Keren et al.
-(2021).
+This scenario is the standard setting for the majority of goal recognition design studies, as highlighted in Table 1 of the survey by Keren et al.
+[15].
 The objective of this set of simulations is to enable comparisons with state-of-the-art methods.
-In our experiments, the initial environment is generated randomly: we first randomly select the number of blocked cells from the range $[0,12]$, followed by the random allocation of the blocked cells’ positions.
+In our experiments, the initial environment is generated randomly: we first randomly select the number of blocked cells from the range $[0,12]$, followed by randomly allocating the blocked cells.
 We also randomly determine the starting grid and two goal grids.
 Environments where the goals are not reachable from the starting grid are filtered out.
-Following this procedure, we randomly generate 200 environments and compare the average performance of our approach with that of baseline methods.
+We randomly generate 200 environments and compare the average performance of our approach with that of baseline methods.
 (a) $6\times 6$ grids
@@
 Figure 3: wcd reduction in a grid world when only blocking modifications are allowed.
-Exhaustive search and Pruned-Reduce are not included in the larger environments of (b) because they cannot complete the computation within an hour for a single environment.
+Exhaustive search and Pruned-Reduce are not included in (b) because they take more than an hour to compute for a single environment.
+(c) Shared budget
+(d) Individual budget
+Figure 4: wcd reduction in settings with two types of modifications.
+We only included greedy baselines as the state-of-the-art baselines, such as Exhaustive search and Pruned-Reduce, are not applicable in these settings.
 We start our experiments with a grid world of size $6\times 6$.
@@
 As illustrated in Figure 3(b), our approach surpasses the greedy baselines in wcd reduction and is 3 times faster than these baselines for large budgets.
+Figure 5: Overcooked-AI Environment. Assuming optimal agent behavior.
+Figure 6: Grid world ($6\times 6$). Incorporating suboptimal agent behavior.
 5.1.2 Settings with flexible budget constraints
@@
 Regarding runtime, our approach is several orders of magnitude faster than the greedy method with true wcd and at least 3 times faster than the greedy method with predicted wcd.
-(a) Shared budget.
-(b) Individual budget.
-Figure 4: wcd reduction in settings with two types of modifications.
 5.1.3 Complex domain and suboptimal agent behavior
@@
 Regarding runtime, our approach is again several orders of magnitude faster than the greedy method with true wcd and at least 3 times faster than the greedy method using predicted wcd.
-Figure 5: Overcooked-AI Environment. Assuming optimal agent behavior.
-Figure 6: Grid world of size $6\times 6$. Incorporating suboptimal agent behavior.
 5.2 Real-World Human-Subject Experiments
@@
 We recruited 200 workers from Amazon Mechanical Turk for this study, paying an average hourly rate of approximately \$16 per worker.
-Figure 7: The interface of our human subject experiment.
-Table 1: The prediction accuracy for human behavior assuming optimal behavior and using data-driven model in Experiment 1.
-Training Accuracy Validation Accuracy Testing Accuracy
-Assuming Optimal Behavior 0.7266 0.6964 0.7131
-Data-Driven Model 0.9189 0.8136 0.8422
 Learning models of human behavior.
-The collected human data were divided into training, validation, and testing sets: the training set included data from 160 workers, comprising approximately 70,000 instances of user decisions, while the validation and testing sets each contained data from 20 workers, amounting to around 8,800 instances of user decisions each.
+The collected human data were divided into training, validation, and testing sets: the training set included data from 160 workers, with approximately 70,000 instances of user decisions, while the validation and testing sets each contained data from 20 workers, with around 8,800 instances of user decisions each.
 We employed a 4-layer Multilayer Perceptron (MLP) model for training the model.
 The input to the model is the current environment layout, and its output predicts the next human action.
-We fine-tuned the hyperparameters, including learning rate, hidden layer size, and L2 penalty, based on the performance on the validation dataset.
+We fine-tuned the hyperparameters based on the performance on the validation dataset.
 We compared the performance of our learned model against a model that assumes optimal agent behavior, which is defined as taking the shortest path to the goal.
@@
 This deviation underscores the importance of incorporating a realistic model of human behavior in goal recognition design, particularly when humans are the decision-makers.
+Table 1: Prediction accuracy of human behavior assuming optimal behavior versus data-driven model.
+Assuming Optimal Behavior Using Data-Driven Model
+Training accuracy 0.7266 0.9189
+Validation accuracy 0.6964 0.8136
+Testing accuracy 0.7131 0.8422
 5.2.2 Experiment 2: Evaluating goal recognition design
@@
   • $\bullet$
-    Initial: No updates to the environment.
+    Original: No updates to the environment.
   • $\bullet$
@@
     Proposed (data-driven): Our proposed approach using predicted wcd from the data-driven model.
-We recruited 200 workers from Amazon Mechanical Turk.
-Each worker was randomly assigned to one of the four treatments specified above, with the distinction between treatments being the environments presented to them.
+We recruited 200 workers from MTurk.
+Each worker was randomly assigned to one of the four treatments above, with the distinction between treatments being the environments presented to them.
 Workers were randomly assigned a goal for each environment and tasked with navigating to reach it.
-After collecting their series of actions aimed at reaching the goal, we analyzed this behavioral data to assess the effectiveness of goal recognition design for each approach.
-Comparing average overlapping path.
-To evaluate the effectiveness of each goal recognition design approach, we first measure the average overlapping path towards each of the two goals.
-Specifically, each recruited worker is exposed to all 30 environments in their assigned treatment and is instructed to reach one of two randomly selected goals.
-For each treatment, we calculate the number of overlapping actions for every pair of workers assigned to different goals within each environment.
-We then compute the average of these numbers, yielding the average overlapping path as our performance measure.
-This metric reflects the difficulty of inferring agent goals for the envioronment and serves as an empirical proxy for wcd.
-With this metric, we vary the initial $k$ portion of actions taken by the workers in our analysis and observe how the average overlapping path changes with the amount of observed actions, denoted by $k$.
-As depicted in Figure 10(c), our approach, when integrated with data-driven models, leads to environments that facilitate the easier identification of the humans’ intended goal.
-Comparing the accuracy of goal inference.
-We next measure the accuracy of inferring human goals based on partial observations, utilizing an off-the-shelf Bayesian inference algorithm [^3] We assume an initial uniform prior distribution.
-As the agent takes actions, we update the posterior belief over goals based on the likelihood probability of the data-driven behavior model prediction.
-.
-Specifically, for each worker, we reveal the first $k$ portion of the worker’s actions to the inference algorithm to infer the worker’s goal.
-This enables us to compute the average inference accuracy.
-The results, as shown in Figure 8(b), demonstrate that our approach leads to environments that are easier for goal recognition.
-(a) Averaging overlapping path (the lower the better.)
+We utilize the collected data to evaluate different goal recognition design approaches.
+(a) Empirical overlapping actions (the lower the better).
 (b) Bayesian goal inference (the higher the better).
@@
 Our approach coupled with data-driven models is shown to generate environments that enable the most effective goal recognition with real human decision makers.
+Comparing empirical overlapping actions.
+To evaluate the effectiveness of each goal recognition design approach, we first measure the empirical overlapping actions towards each of the two goals.
+Specifically, each recruited worker is exposed to all 30 environments in their assigned treatment and is instructed to reach one of two randomly selected goals.
+For each treatment, we calculate the number of overlapping actions for every pair of workers assigned to different goals within each environment.
+Therefore, for each treatment, we obtain a distribution of the number of overlapping actions.
+This metric reflects the difficulty of inferring agent goals for the environment and serves as an empirical proxy for wcd.
+Figure 7(a) presents the percentiles of the number of overlapping actions for each treatment.
+In particular, in the lowest 25%, the environments by our approaches have no overlapping actions, indicating that the environment makes it easy to infer the goal of the agents from observations.
+Generally, our approach, when integrated with data-driven models, leads to environments with fewer number of overlapping actions and therefore facilitate the easier identification of the humans’ intended goal.
+Comparing the accuracy of goal inference.
+Instead of using wcd or proxy of wcd for evaluations, we next directly measure whether we can indeed infer the human goals based on partial observations of their actions, utilizing an off-the-shelf Bayesian inference algorithm.
+[^3] We assume an initial uniform prior distribution.
+As the agent takes actions, we update the posterior belief over goals based on the likelihood probability of the data-driven behavior model prediction.
+Specifically, for each worker, we reveal the first $k$ portion of the worker’s actions to the inference algorithm to infer the worker’s goal.
+This enables us to compute the average inference accuracy.
+The results in Figure 7(b) demonstrate that our approach leads to environments that are easier for goal recognition.
 6 Conclusion
-Effective collaboration between humans and AI hinges on the AI’s ability to infer intentions and goals of humans.
-In this study, we work on the problem of goal recognition design, aiming to update the decision-making environments to make it easier for the AI agents to perform goal recognition.
+Effective human-AI collaboration hinges on the AI’s ability to infer the goals of humans.
+In this study, we work on the problem of goal recognition design, updating the decision-making environments to make it easier for the AI agents to perform goal recognition.
 By addressing the key limitations in the existing literature, notably the computational demand and the assumption of optimal agent behavior, we have developed a data-driven optimization framework that is both efficient and scalable.
-Moreover, our approach is capable of integrating general models of various agent behavior, moving beyond the traditional assumption of optimality and making it more applicable to real-world scenarios involving human agents.
-Future work of our study includes extending our framework to more complex environments, particularly those that are dynamic and partially observable.
-Given our framework is adaptable to different forms of agent behavior, understanding how different patterns of agent suboptimality impacts the difficulty of goal recognition could lead to useful insights for enabling better human-AI collaboration.
+Through simulations and human-subject experiments, we show that our approach outperforms state-of-the-art approaches in standard settings, applies to settings that existing approaches cannot address, and leads to environments easier for goal recognition with real-world human decision makers.
+Societal impact.
+Our approach is capable of integrating general models of various agent behavior, making it more applicable to real-world scenarios involving human agents.
+However, with more accurate human models, it also opens up the concerns of privacy and potential mis-use of the techniques for unethical surveillance.
+While our work does not directly facilitate such misuse, the broader field of goal recognition research could potentially be exploited.
+Limitations and future work.
+While our work has extended the literature in goal recognition design to address suboptimal agent behavior, more flexible budget constraints, and more complicated environments, there are still several limitations.
+In particular, we have assumed static and full observability of the environment.
+Extending our framework to deal with more complex environments, particularly those that are dynamic and partially observable would be an interesting and important future work.
+Moreover, we have assumed that user behavior remains static, so we can learn a human model from their historical behavior.
+Finally, given our framework is adaptable to different forms of agent behavior, understanding how different patterns of agent suboptimality impacts the difficulty of goal recognition could lead to useful insights for enabling better human-AI collaboration.
+Acknowledgements
+This work is supported in part by J.P. Morgan Faculty Research Award and a Global Incubator Seed Grant from McDonnell International Scholars Academy.
 Appendix A Experiment Details and Additional Results
@@
 In the standard setup with a small grid size, all methods except exhaustive complete within less than half of a second but only our approach scales with a larger grid size and other more complex configurations.
+All our experiments are conducted on a cluster of 40 CPU cores (Intel Xeon Gold 6148 CPU @@@@2.40GHz), 1 GPU (NVIDIA Tesla V100 SXM2 32GB), and a maximal memory of 80GB.
 Appendix B Additional Experiment Results
@@
 Figure 10: wcd reduction in smaller ($6\times 6$) gridworld with optimal agent behavior.
-Appendix C Simulation Details
+B.1 Definition of Budget
+In our evaluation, we compare the wcd reduction relative to the budget allocated for modifications.
+In the grid-world domain, the budget represents the number of changes made, which are limited to two types: blocking or unblocking cells.
+In contrast, the Overcooked-AI domain allows for a richer set of modifications due to its complexity.
+Here, modifications involve changing the positions of objects.
+A valid environment must include all specified objects, as detailed in Section 4.1.2.
+The budget in this domain is quantified as the total Manhattan distance moved by the objects between the original and modified environments.
+Appendix C Details of Simulations
 To generate our training and evaluation datasets, we randomly generated environments and kept environments that are valid, e.g., the goals are reachable, and that the objects don’t overlap.
 Below we provide more details for environment generation for specified grid sizes for both overcooked and grid-world domains.
-C.0.1 Gridworld
+C.1 Gridworld
 In a grid world, a valid environment includes a starting position, blocked cells, and two goal positions.
@@
 The two goal states assigned a large reward that is 10 times the largest subgoal reward.
-C.0.2 Overcooked-AI
+C.2 Overcooked-AI
 In Overcooked-AI, a valid environment includes one pot, one tomato source, one onion source, one dish source, one serving point, no open spaces at the border, and any number of open spaces and blocked cells.
@@
 The task interface is shown in Figure 11.
-Figure 12: Human-Subject experiment interface.
+Note that while each environment has two goals, we only show one goal (the goal of the worker) to the worker in our interface to simplify the presentations.
+The second goal is shown as a blocked cell in the interface, i.e the worker only navigates to the shown goal.
+Figure 11: The interface of our human subject experiment.
 We have recruited 200 workers from Amazon Mechanical Turk in total, and Table 2 contains the demographic information of the workers.

### v3

@@
-Data-Driven Goal Recognition Design for General Behavioral Agents
+Goal Recognition Design for General Behavioral Agents using Machine Learning
 Abstract
-Goal recognition design aims to make limited modifications to decision-making environments with the goal of making it easier to infer the goals of agents acting within those environments.
+Goal recognition design (GRD) aims to make limited modifications to decision-making environments to make it easier to infer the goals of agents acting within those environments.
 Although various research efforts have been made in goal recognition design, existing approaches are computationally demanding and often assume that agents are (near-)optimal in their decision-making.
-To address these limitations, we introduce a data-driven approach to goal recognition design that can account for agents with general behavioral models.
+To address these limitations, we leverage machine learning methods for goal recognition design that can both improve run-time efficiency and account for agents with general behavioral models.
 Following existing literature, we use worst-case distinctiveness (wcd) as a measure of the difficulty in inferring the goal of an agent in a decision-making environment.
@@
 We then propose a gradient-based optimization framework that accommodates various constraints to optimize decision-making environments for enhanced goal recognition.
-Through extensive simulations, we demonstrate that our approach outperforms existing methods in reducing wcd and enhancing runtime efficiency in conventional setup.
-Morevoer, our approach also adapts to settings in which existing approaches do not apply, such as those involving flexible budget constraints, more complex environments, and suboptimal agent behavior.
-Finally, we have conducted human-subject experiments which confirm that our method can create environments that facilitate efficient goal recognition from real-world human decision-makers.
+Through extensive simulations, we demonstrate that our approach outperforms existing methods in reducing wcd and enhances runtime efficiency.
+Moreover, our approach also adapts to settings in which existing approaches do not apply, such as those involving flexible budget constraints, more complex environments, and suboptimal agent behavior.
+Finally, we conducted human-subject experiments that demonstrate that our method creates environments that facilitate efficient goal recognition from human decision-makers.
 1 Introduction
-With the rapid advancement of artificial intelligence (AI), there has been a surge in interest in human-AI collaboration, aiming to synergize human and AI capabilities across various domains such as gaming, e-commerce, healthcare, and workflow productivity.
+With the rapid advancement of artificial intelligence (AI), there has been a surge in interest in human-AI collaboration, aiming to synergize human and AI capabilities across various domains, such as gaming, e-commerce, healthcare, and workflow productivity.
 Designing AI agents to work alongside humans requires these agents to understand and infer human goals and intentions.
-While there has been abundant research in goal recognition [23, 24] that aims to infer human goals by observing their actions, this work focuses on the goal recognition design problem [12], where one needs to identify how to modify decision-making environments to enable better goal recognition.
-Our work is built on the goal recognition design problem formulated by Keren et al.
-[12].
+While there has been abundant research in goal recognition (Ramírez & Geffner, 2010; Sukthankar et al., 2014), aiming to infer human goals by observing their actions, this work focuses on the goal recognition design problem (Keren et al., 2014), which seeks to identify how to modify decision-making environments to make it easier to infer human goals.
+Our work is built on the GRD problem formulated by Keren et al.
+(2014).
 They proposed the worst-case distinctiveness (wcd) metric, defined as the maximum number of decisions an agent can make without revealing its goal, to measure the difficulty of inferring the agent’s goal.
-They then aimed to modify the decision-making environment, through removing allowable actions from states, to optimize this measure.
-Since the introduction of this work, there have been several follow-up works that extend the formulation to deal with different settings, such as stochastic settings [28, 29] and partial observability [14].
+They then aimed to modify the decision-making environment, by removing allowable actions from states, to optimize this measure.
+Since the introduction of this work, there have been several follow-up works to deal with different settings, such as stochastic settings (Wayllace et al., 2016; 2017) and partial observability (Keren et al., 2016).
 More discussion can be found in the survey by Keren et al.
-[15].
-While there has been significant progress in goal recognition design, there are two main limitations in the literature.
-First, current approaches require evaluating the difficulty of goal recognition, i.e., worst-case distinctiveness (wcd), for a large number of potential modifications to the decision-making environment.
-Since each evaluation of wcd requires solving the optimal policy multiple times to each of the goals, this process is computationally demanding and limits the scalability of these approaches, especially in virtual domains where the number of possible environment configurations is large.
-Second, decision-making agents in goal recognition design are often assumed to be optimal.
-While there are efforts addressing suboptimal agents [13, 27], these primarily focus on settings where there is a limited number of deviations from optimal decision-making.
-This assumption of optimality is limiting in scenarios involving human agents, who are known to often systematically deviate from optimal decision-making due to cognitive and informational constraints [10, 19].
-To address these limitations, we propose a framework for data-driven goal recognition design with general agent behavior.
-To relax the optimality assumption, we explicitly incorporate models of agent behavior into the optimization framework to better represent behavioral agents.
+(2021).
+While there has been significant progress in GRD, there are two main limitations in the literature.
+First, existing approaches often assume that decision-making agents are optimal or near-optimal decision makers (Keren et al., 2015; Wayllace & Yeoh, 2022).
+This assumption is unrealistic with human decision makers, who are known to often systematically deviate from optimal decision-making due to cognitive and informational constraints (Kahneman, 2003; O’Donoghue & Rabin, 1999).
+Second, existing approaches require evaluating the difficulty of goal recognition, i.e., worst-case distinctiveness (wcd), for a large number of modifications to the environment.
+This computational requirement limits the scalability of these approaches, especially for virtual environments when there might be frequent updates of the environment, e.g., e-commerce platforms might want to frequently adjust the website design to enhance user intent inference.
+We propose a framework for GRD with general agent behavior to address these limitations.
+To relax the optimality assumption, we explicitly incorporate models of agent behavior into the optimization framework.
 To tackle the computational challenges, our approach leverages data-driven methods for goal recognition design.
-The core idea involves building a machine learning oracle that predicts the difficulty of goal recognition (e.g., wcd) given a decision-making environment and an agent’s behavioral model.
-This oracle is trained on datasets generated from simulations, allowing for the evaluation of wcd in any given environment and agent model.
-Such an approach significantly accelerates the evaluation of wcd during run-time optimization.
-Once the machine learning oracle is established, we apply the general gradient-based optimization approach to minimize wcd, employing Lagrangian relaxation to manage constraints on environment modifications.
-This allows us to address more general forms of objectives and constraints that existing approaches in the literature cannot address.
+The core idea is to build a machine learning oracle that predicts the difficulty of goal recognition (e.g., wcd) given a decision-making environment and an agent’s behavioral model.
+This oracle is trained on datasets generated from simulations, enabling efficient evaluation of wcd for any given environment and agent model.
+Such an approach significantly accelerates the evaluation process during optimization.
+Once the machine learning oracle is established, we apply a general gradient-based optimization method to minimize wcd, using Lagrangian relaxation to handle constraints on environment modifications.
+Beyond addressing the two main limitations, this optimization procedure also supports more general forms of objectives and constraints that existing approaches in the literature cannot accommodate.
 To evaluate our framework, we conducted a series of simulations and human-subject experiments.
-We start with simulations in the standard setup, with the grid world environment and optimal agent assumption.
-We show that our approach outperforms existing baselines in reducing worst-case distinctiveness (wcd) and demonstrates considerably better run-time efficiency.
-We then conducted additional simulations to showcase that our approach can generalize to settings beyond the capabilities of existing approaches in the literature.
-These include scenarios involving flexible budget constraints, more complex environments, and suboptimal agent behavior.
+We start with simulations in the standard setup in the literature, with the grid world environment and optimal agent assumption.
+We show that our approach outperforms existing baselines in reducing worst-case distinctiveness (wcd) and achieves considerably better run-time efficiency.
+We then conducted additional simulations to demonstrate that our approach can generalize to settings that existing methods cannot address, including scenarios involving flexible budget constraints, more complex environments, and suboptimal agent behavior.
 Lastly, we have conducted human-subject studies demonstrating that our method can be leveraged to design environments that facilitate efficient goal recognition from real-world human decision-makers.
@@
   • $\bullet$
-    We propose a data-driven optimization framework for goal recognition design.
-This framework comprises a predictive model that estimates the wcd for a given environment and a model of agent behavior.
-We then utilize a gradient-based optimization method for goal recognition design.
-The framework is flexible to accommodate various environments, design spaces, and agent behavior.
+    We propose a data-driven optimization framework that accommodates general agent behavior and improves the efficiency of goal recognition design, addressing two main limitations in the existing literature while enabling a more flexible optimization procedure.
+The framework consists of a predictive model that estimates the wcd for a given environment and a model of agent behavior.
+We then apply a gradient-based optimization method to perform goal recognition design.
+The framework is runtime efficient, capable of incorporating general agent behavior, and adaptable to different design spaces and constraints.
   • $\bullet$
-    Through extensive simulations, we show that our framework not only outperforms existing approaches in goal recognition design within standard settings but also adapts to scenarios that existing approaches cannot handle, including general optimization criteria, complex environments, and suboptimal agent behavior.
+    Through extensive simulations, we show that our framework not only outperforms existing approaches in goal recognition design, both in reducing worst-case distinctiveness wcd and improving run-time efficiency in standard settings, but also adapts to scenarios that existing methods cannot handle, including general optimization criteria, complex environments, and suboptimal agent behavior.
   • $\bullet$
-    Through human-subject experiments, we demonstrate that our approach can adapt to agent models trained on human behavior.
-Furthermore, our framework can create environments that enable more effective goal recognition with real-world humans.
-To the best of our knowledge, our work is the first to utilize human-subject experiments to evaluate the environment design for goal recognition.
+    Through human-subject experiments, we demonstrate that our approach adapts to settings in which decision-making agents are real-world humans.
+To the best of our knowledge, our work is the first to evaluate the environment design for goal recognition with human-subject experiments.
 2 Related Work
-Our work contributes to the expanding field of human-AI collaboration.
-Recent research has indicated that optimizing AI alone is insufficient for maximizing the performance of human-AI teams [1, 2].
-To develop truly collaborative AI agents, it is crucial to equip them with the ability to comprehend and predict the intentions and goals of their human counterparts.
-This challenge lies at the heart of goal recognition research [11, 23, 9, 24, 20, 17].
-Our work specifically focuses on goal recognition design, an extension of goal recognition that includes modifying environments to better facilitate the process of recognizing goals.
-Goal recognition design was formulated by Keren et al.
-[12].
-Since this seminal work, numerous research efforts have expanded the concept to accommodate stochastic environments [29, 27], different levels of observability [14, 30], and a variety of design spaces [18].
-The studies most closely aligned with our approach focus on suboptimal agents [13, 27].
-However, these studies characterize suboptimality by limiting deviations from the optimal policy, a method that may not adequately represent behavior of human agents, who frequently deviate systematically from optimal decision-making.
-Additionally, most existing work requires evaluating the difficulty of goal recognition in run time for numerous environmental modifications, which constrains the scalability of these methods.
-Our work sets itself apart by broadening the goal recognition design question to incorporate general agent behavioral models and by implementing a data-driven optimization approach.
-From a technical standpoint, we adopt a data-driven optimization approach for goal recognition design.
-The use of data-driven tools in optimization has gained increasing prominence in the field of mechanism design, as evidenced by a range of recent studies [5, 7, 8, 4, 16, 22, 21, 3, 32].
-Moreover, our work closely aligns with the recent studies that explicitly encoded human behavior and responses in the design of computational and learning systems [25, 26, 31, 32, 6].
-While our work shares a similar motivation, our approach and problem formulation differ from theirs.
+Our work contributes to the field of human–AI collaboration.
+Recent research shows that optimizing AI alone is insufficient to maximize the performance of human–AI teams (Bansal et al., 2021; Carroll et al., 2019; Yu et al., 2024), motivating efforts to understand and model human behavior in interactions with AI systems (Choudhury et al., 2019; Shah et al., 2019; Kwon et al., 2020; Chan et al., 2021; Reddy et al., 2021; Narayanan et al., 2022a; b; 2023; Treiman et al., 2023; 2024; 2025).
+To build truly collaborative AI agents, they must also anticipate the intentions and goals of their human counterparts.
+This challenge lies at the core of goal recognition research (Kautz, 1991; Ramírez & Geffner, 2010; Hong, 2001; Sukthankar et al., 2014; Pereira et al., 2017; Masters et al., 2021).
+Our work focuses on goal recognition design, an extension of goal recognition that incorporates environment modifications to make goals easier to recognize.
+Goal recognition design (GRD) was formulated by Keren et al.
+(2014).
+Since this seminal work, numerous research efforts have extended to accommodate stochastic environments (Wayllace et al., 2017; Wayllace & Yeoh, 2022), varying levels of observability (Keren et al., 2016; Wayllace et al., 2020), and diverse design spaces (Mirsky et al., 2019).
+The studies most closely aligned with our approach focus on suboptimal agents (Keren et al., 2015; Wayllace & Yeoh, 2022).
+However, these studies characterize suboptimality by limiting deviations from the optimal policy, a method that may not adequately capture the behavior of human agents, who frequently deviate systematically from optimal decision-making.
+Moreover, most existing work requires evaluating goal recognition difficulty at run time across a large number of environmental modifications, which significantly limits scalability.
+Our work distinguishes itself by broadening GRD to incorporate general models of agent behavior and by implementing a data-driven optimization approach.
+There has been effort to accelerate GRD by using heuristic approaches that avoid the prohibitively large search space of possible modifications.
+In their seminal work, Keren et al.
+(2014) proposed the pruned-reduce method, which restricts modifications to action removal and leverages a key insight: blocking actions do not introduce new goal-directed paths.
+This alleviates the computational burden of exhaustive search.
+More recent work by Pozanco et al.
+(2024) accelerates GRD by using a plan library from a top-quality planner (Katz et al., 2020) to guide initial heuristic exploration.
+Similarly, Au (2024) expanded the design space with block-level modifications—simultaneous, interconnected changes—to improve efficiency.
+While these approaches make notable strides, they still rely on exhaustive search and repeated computation of complex metrics, which remains the primary computational bottleneck.
+In contrast, our method adopts a data-driven optimization framework.
+Rather than evaluating the metric (wcd) at each step of the optimization process, we train predictive models to approximate it directly, enabling rapid evaluation and comparison of design candidates.
+Moreover, the differentiability of the predictive model allows for gradient-based optimization in place of exhaustive search.
+Together, these advances significantly reduce computational overhead and open new possibilities for scaling GRD to more complex settings.
+Our methodology aligns with recent advances in data-driven optimization for mechanism design (Dütting et al., 2019; Golowich et al., 2018; Curry et al., 2020; Kuo et al., 2020; Rahme et al., 2020; Peri et al., 2021; Cornelisse et al., 2022; Yu et al., 2023), which leverage machine learning to accelerate otherwise intractable tasks.
+It also aligns with recent research that incorporates human models into learning and optimization processes (Ho et al., 2016; Evans et al., 2016; Shah et al., 2019; Tang & Ho, 2019; 2021; Carroll et al., 2019; Tang et al., 2021a; b; Yu & Ho, 2022; Feng et al., 2024; Kasumba et al., 2025).
+While we share the high-level motivation, our problem formulation and methodological choices offer novel contributions and advance the study of goal recognition design through a principled integration of predictive modeling and optimization.
 3 Problem Formulation and Methods
 3.1 Problem Setting
-We formulate the decision-making environment and behavioral agent models. We also explain the commonly-used metrics for evaluating goal recognition difficulty.
 Decision-making environment.
@@
 The agent is conceptualized as a planner $H:W\rightarrow\Pi$, where the input is an environment $w\in W$, and the output is a policy $\pi\in\Pi$.
-To illustrate our formulation, consider a agent parameterized by a time-variant discounting function $d(t)$.
+To illustrate our formulation, consider an agent parameterized by a time-variant discounting function $d(t)$.
 The standard optimal agent model corresponds to a fixed discounting factor $\gamma\in(0,1]$ with $d(t)=\gamma^{t}$.
-We can also represent an agent with present bias [19] by adopting a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$, where $k>0$.
-We would like to highlight that our approach not only accounts for standard analytical closed-form expression of agent behavior.
-It can also accounts for scenarios where an agent policy $\pi$ is a machine learning model, i.e., a neural network trained on human behavioral data.
+We can also represent an agent with present bias (O’Donoghue & Rabin, 1999) by adopting a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$, where $k>0$.
+Note that our approach not only accounts for standard analytical closed-form expression of agent behavior.
+It can also account for scenarios where an agent policy $\pi$ is a machine learning model, i.e., a neural network trained on human behavioral data.
 Worst-case distinctiveness (wcd).
-To evaluate the difficulty of goal recognition, we adhere to standard literature and focus on the measure of worst-case distinctiveness (wcd[12], defined as the maximum number of steps an agent can take before revealing its goal.
-To compute the wcd for a given suboptimal agent $h\in H$ in an environment $w\in W$, we evaluate the path for the agent to each goal and compute the number of actions from the initial state that are identical for every goal.
-3.2 Goal Recognition Design with General Behavioral Agents
-We formalize the goal recognition design problem with general behavioral agents.
+To evaluate the difficulty of goal recognition, we follow the standard literature and focus on worst-case distinctiveness (wcd(Keren et al., 2014), defined as the maximum number of steps an agent can take that are consistent with multiple goals before its true goal becomes distinguishable.
+In other words, it measures the number of initial steps that overlap across trajectories toward different goals.
+To compute the wcd for a given agent $h\in H$ in an environment $w\in W$, we evaluate the path for the agent to each goal and compute the number of actions from the initial state that are identical for each goal.
+While we focus on wcd in this work, our method is agnostic to the specific metric used, as long as the metric is computable, a predictive model can be trained to approximate it and enable efficient evaluation.
+3.2 Goal Recognition Design (GRD) Formulation
+We formalize the GRD problem with general behavioral agents.
 Given an environment $w$ and an agent with a behavior model $h$, we denote the worst-case distinctiveness of environment $w$ for agent $h$ as $\textit{wcd}(w,h)$.
 Each type of modification $1\leq i\leq N$ will incur a cost $c_{i}(w,w^{\prime})$ that must fall in budget constraint $B_{i}$.
-The objective of the goal recognition design problem is to alter the environment from $w$ to $w^{\prime}$ in a way that minimizes $\textit{wcd}(w^{\prime},h)$, while satisfying the constraint that the cost of the modifications does not exceed the budget.
+The objective of the GRD problem is to alter the environment from $w$ to $w^{\prime}$ in a way that minimizes $\textit{wcd}(w^{\prime},h)$, while satisfying the constraint that the cost of the modifications does not exceed the budget.
 $\underset{w^{\prime}}{\text{minimize}}$ $\textit{wcd}(w^{\prime},h)$ (1)
@@
 3.3 Our Proposed Method
-Existing approaches in goal recognition design require evaluating wcd for a large number of potential modifications to the environments to identify the optimal modifications.
-Since evaluating wcd requires evaluating the agent policy for multiple goals, it presents significant computational overhead and limits the scalability of these approaches.
-To overcome this challenge, we propose leveraging machine learning to expedite run-time computations.
-The main idea, as summarized in Figure 1, is to first train a machine learning model that predicts wcd for any given pair of decision-making environment $w$ and agent behavioral model $h$.
-After obtaining this machine learning model and utilizing its differentiable property, we develop an optimization framework that applies gradient-based optimization methods to the Lagrangian relaxation of the constrained optimization formulation in (1).
-The main benefits of our approach compared with existing methods include: 1) the ability to incorporate different agent models $h$, 2) flexibility for various forms of optimization objectives and constraints, and 3) run-time efficiency.
+Existing approaches to GRD require evaluating wcd for a large number of potential environment modifications in order to identify the optimal ones.
+Since computing wcd involves evaluating the agent’s policy across multiple goals, it introduces significant computational overhead and relies on strong assumptions about agent behavior.
+To address this challenge, we propose leveraging machine learning to expedite runtime computation and explicitly account for general agent models.
+The main idea, illustrated in Figure 1, is to first train a machine learning model that predicts wcd for any given pair of decision-making environment $w$ and agent behavioral model $h$.
+Once this predictive model is trained, we exploit its differentiability to develop an optimization framework that applies gradient-based methods to the Lagrangian relaxation of the constrained optimization problem defined in Equation 1.
+This approach not only addresses the two key limitations—computational challenges and restrictive assumptions about agent behavior—but also provides flexibility for accommodating various forms of optimization objectives and constraints.
 (a) Predictive model for wcd.
 (b) Optimization framework.
-Figure 1: We first train a predictive model to predict wcd from simulated data.
-We then perform gradient-based optimization that leverages the predictive model to identify environment modifications that minimize wcd with a given model of agent behavior.
-(a) Grid world.
-(b) Overcooked-AI.
-Figure 2: The example showcases two benchmark environments.
-The first, displayed on the left, is a grid world environment.
-The agent starts at a position marked ’S’ and aims to reach one of the goal positions labeled ’G’.
-The agent must navigate through the grid, avoiding blocked cells marked with ’x’.
-The second environment, illustrated on the right, is an Overcooked-AI setting.
-In this scenario, the agent’s objective is to pick up ingredients and complete the cooking of their target recipe, which constitutes their goal.
+Figure 1: We first train a wcd predictor from simulated data.
+We then perform gradient-based optimization that leverages the predictor to identify environment modifications that minimize wcd with a given agent behavior model.
 Predictive model for wcd.
 To build the predictive model for wcd, we curate a training dataset through simulations.
-For a given environment $w$ and agent behavioral model $h$, we can obtain $\textit{wcd}(w,h)$ by solving for the agent’s actions towards different goals.
+For an environment $w$ and agent behavioral model $h$, we can obtain $\textit{wcd}(w,h)$ by solving for the agent’s actions towards different goals.
 After collecting a training dataset, we train the predictive model using a convolutional neural network.
-Implementation details are included in the appendix.
+The implementation details are in Section 5.1.1 and the appendix.
 Optimization procedure.
 After obtaining the predictive model, we develop a gradient-based optimization framework.
-This framework strictly generalizes the existing literature in goal recognition design where the space of modifications are limited (e.g., usually limits to blocking an action in MDP) [15].
-The first step involves transforming the constrained optimization problem in (1) into an unconstrained optimization problem using Lagrangian relaxation:
+This framework generalizes the existing literature in GRD where the space of modifications is limited (e.g., usually limits to blocking an action in MDP) (Keren et al., 2021).
+The first step of the optimization procedure involves transforming the constrained optimization problem in equation 1 into an unconstrained optimization problem using Lagrangian relaxation:
 $\mathcal{L}=\textit{wcd}(w^{\prime},h)+\sum_{i}\lambda_{i}(c_{i}(w,w^{\prime})-B_{i}).$
-We then perform gradient descent on the relaxed Lagrangian.
-As environment modifications are often discrete (e.g., whether to block a cell in the grid world), we apply a discrete gradient descent procedure.
-Specifically, at each step of gradient descent, we obtain a gradient, which is a vector indicating the suggested change magnitude for each element (such as a cell in the grid world).
-We then select the element with the highest gradient value and make the corresponding change.
-Note that some suggested modifications may not be valid; for instance, we cannot block a cell that is already blocked in the grid world.
+We then perform gradient descent on the relaxed Lagrangian with respect to $w$.
+As environment modifications are often discrete (e.g., to block a cell in the grid world), we apply a discrete gradient descent procedure.
+Specifically, at each descent step, we obtain a gradient, which is a vector indicating the suggested change magnitude for each element (such as a cell in the grid world).
+We then select the element with the gradient value of the largest magnitude and make the corresponding change.
+Note that some suggested modifications may be invalid; for instance, we cannot block an already blocked cell in the grid world.
 In such cases, we proceed to the element with the next highest gradient value, continuing this process until a valid modification is made.
 This modification procedure is repeated until the gradient descent converges.
-Note that with this Lagrangian relaxation, we cannot directly set the budget, however, based on duality, a larger Lagrangian multiplier $\lambda$ corresponds to a smaller budget $B$, in the original constrained optimization formulation.
-In our experiments, instead of selecting different budgets directly, we choose varying Lagrangian multipliers and record the realized costs of the modifications.
+Note that with this Lagrangian relaxation, while we cannot directly set the budget, based on duality, a larger Lagrangian multiplier $\lambda$ corresponds to a smaller budget $B$ in the original constrained optimization formulation.
+In practice, designer can choose to vary the size of the Lagrangian multiplier to modulate the budget.
+In our experiments, we choose varying Lagrangian multipliers and record the realized costs of the modifications.
 4 Experiment Setup
@@
 We utilize two benchmark domains.
-The first is the standard basic grid world environment, commonly used in goal recognition design literature.
-The second is the Overcooked-AI environment [2], an complex environment with a richer set of environment modification.
+The first is the standard basic grid world environment, commonly used in the GRD literature.
+The second is the Overcooked-AI environment (Carroll et al., 2019), a more complex environment with a richer set of environment modifications.
 This environment is particularly relevant to the downstream implications of our work, namely in supporting human-AI collaboration.
+(a) Grid world.
+(b) Overcooked-AI.
+Figure 2: The benchmark environments.
+The left one is a grid world.
+The agent starts at a position marked ’S’ and aims to reach one of the goal positions labeled ’G’.
+The agent must navigate through the grid, avoiding blocked cells marked with ’x’.
+The right one is an Overcooked-AI setting, where the agent’s objective is to pick up ingredients and complete their target recipe, which constitutes their goal.
 4.1.1 Benchmark Domain: Grid World
 In the grid world domain, agents navigate a grid with several potential goals.
-Take, for example, a grid world environment as illustrated in Figure 2(a).
-In this environment, an optimal agent starts at point ’S’ and aims for one of the goals.
-Spaces marked ’X’ are blocked, barring the agent’s passage.
-In this particular environment, the worst-case distinctiveness (wcd) is 0, as an optimal agent, following the shortest path to either goal, would reveal its intended goal on the first move due to non-overlapping paths.
-Our experiments are primarily focused on grid world environments with two goals for simplicity.
-However, our approach is applicable to scenarios with more than two goals.
+Consider an example in Figure 2(a): an agent starts at point ’S’ and aims for one of the goals marked ’G’.
+Spaces marked ’X’ are blocked.
+In this example, the worst-case distinctiveness (wcd) for an optimal agent is 0, as the shortest paths to the two goals do not overlap, meaning the agent’s intended goal is revealed on the very first move.
+Our experiments primarily focus on grid world environments with two goals for simplicity.
+However, our approach extends naturally to settings with more than two goals by computing the relevant GRD metric and training a corresponding predictive oracle.
 Design space of environment modifications.
-In the context of goal recognition design in grid world, the space environment modification is often limited to adding blocks to spaces [12], also called action removal in the literature.
-In our work, we broaden the design space for modifications to also consider the unblocking of existing blocked spaces as potential modifications.
+In the context of GRD in grid world, the space of environment modification is often limited to adding blocks to spaces in the literature (Keren et al., 2014), also called action removal.
+In our work, we broaden the design space to also consider the unblocking of existing blocked spaces as potential modifications.
 4.1.2 Benchmark Domain: Overcooked-AI
-We also conduct our evaluations on a more complex domain: Overcooked-AI [^1] https://github.com/HumanCompatibleAI/overcooked_ai.
-This environment is based on the popular game Overcooked.
-In each game, players (agents) collaborate to prepare and deliver specific recipes.
-Since the goal of our approach is to enable efficient inference of agent goals, we focus on the special case with a single agent.
-The environment is represented as a grid (see Figure 2(b)) with each cell specifying the object that is placed at the cell.
-The objects may be a counter, a tomato, an onion, a pot, a dish, a serving point, or an empty space.
-The agent can only occupy empty spaces in the environment and cannot step on any other objects in the environment.
-The agent can carry the movable objects, i.e.
-onion, tomato, and dish, and drop them on any of the other non-space objects (counter, serving point, or pot).
-To navigate in the environment, the agent can move left, right, north, or south and maintain an orientation that is consistent with the last movement direction.
+We also conduct our evaluations in a more complex domain: Overcooked-AI.
+[^1] https://github.com/HumanCompatibleAI/overcooked_ai This environment is based on the popular game Overcooked, where players (agents) collaborate to prepare and deliver specific recipes.
+Since the goal of our approach is to enable efficient inference of agent goals, we focus on the simplified setting with a single agent.
+The environment is represented as a grid (see Figure 2(b)), with each cell indicating the object located there.
+Objects may include a counter, tomato, onion, pot, dish, serving point, or empty space.
+The agent can only occupy empty spaces and cannot step on any other object.
+It can carry movable items, i.e., onions, tomatoes, and dishes, and place them on other non-space objects such as counters, serving points, or pots.
+To navigate the environment, the agent can move left, right, up, or down and maintains an orientation aligned with its most recent movement direction.
 Goal recognition in Overcooked-AI.
-The goal of the agent in Overcooked-AI is defined by the recipe it needs to complete.
-For instance, to prepare a soup with one onion and two tomatoes, the agent must pick up the ingredients, place them in the pot, and cook them.
-In the context of goal recognition, the objective is to deduce which recipe the agent is preparing.
-For the sake of simplicity in our experiments, we focus on scenarios with only two possible goals or recipes.
+The agent’s goal in Overcooked-AI is defined by the recipe it needs to complete.
+For example, to prepare a soup with one onion and two tomatoes, the agent must collect the ingredients, place them in a pot, and cook them.
+In the context of goal recognition, the objective is to infer which recipe the agent is pursuing based on its behavior.
+For simplicity, our experiments focus on scenarios with only two possible goals or recipes.
 Design space of environment modifications.
 A primary challenge in the Overcooked-AI environment is its considerably larger design space for environment modifications.
-The design space includes the changing position of any object within the environment, subject to the modification being valid.
-This means that the change cannot result in a new design where any of the objects is unreachable by the agent.
-The objective of conducting experiments in this domain is to assess whether our approach can be used to address more complex domains.
+The design space includes changing the position of any object in the environment, subject to the constraint that the modification remains valid.
+This means that a change cannot result in a new environment design where any of the objects is unreachable by the agent.
+The objective of conducting experiments in this domain is to assess whether our approach can address more complex domains.
 4.2 Baselines
-We now describe the baseline approaches we use to compare against our approach.
+We compare our approach against four baselines that collectively represent the major algorithmic directions in GRD: (Optimal) exhaustive search, heuristic-enhanced search, and greedy-based approaches.
+For heuristic-enhanced search, we implement the Pruned-Reduce method proposed by Keren et al.
+(2014) as a representative example.
+For greedy-based approaches, we implement methods that use either the true or predicted wcd values to prioritize locally beneficial modifications.
   • $\bullet$
@@
 It is guaranteed to achieve minimum wcd.
-However, given the computation overhead, this approach is not applicable in most situations.
+However, given the computational overhead, this approach is not applicable in most situations.
   • $\bullet$
-    Pruned-Reduce [12]: This baseline is specifically designed for settings where modifications are limited to action removal such as blocking a cell in grid world.
+    Pruned-Reduce (Keren et al., 2014): This baseline is specifically designed for settings where modifications are limited to action removal such as blocking a cell in grid world.
 It speeds up the exhaustive search and retains the optimal property.
@@
     Parameterized suboptimal agent behavior.
-We also consider an generalized behavior model parameterized by $d(t)$.
+We consider a generalized behavior model parameterized by $d(t)$.
 The agent’s objective is to optimize a time discounted reward with the discounting factor for $t$ steps in the future being $d(t)$.
 The standard optimal agent model corresponds to a fixed discounting factor $\gamma\in(0,1]$ with $d(t)=\gamma^{t}$.
-We can also represent an agent with present bias [19] by adopting a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$, where $k>0$.
+We can also represent an agent with present bias (O’Donoghue & Rabin, 1999) by adopting a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$, where $k>0$.
   • $\bullet$
@@
 5.1 Simulations
-In our simulations, we first examine how our approach compares with existing methods in standard setups found in the literature.
-We then showcase the generalizability of our approach by extending it to scenarios that go beyond the setups studied in the literature, including scenarios with more dynamic budget constraints, more complicated environments, and suboptimal agent behavior.
-5.1.1 Standard setting
-We first focus on settings within the grid world domain, where modifications are limited to blocking cells and agent behavior is assumed to be optimal.
-This scenario is the standard setting for the majority of goal recognition design studies, as highlighted in Table 1 of the survey by Keren et al.
-[15].
+In our simulations, we first evaluate how our approach compares with existing methods in standard setups commonly found in the literature.
+We aim to show that our method matches or exceeds existing approaches in reducing wcd while achieving significantly better time efficiency.
+We then demonstrate the generalizability of our approach by applying it to scenarios that state-of-the-art methods cannot address, including those with dynamic budget constraints, more complex environments, and suboptimal agent behavior.
+Standard setting.
+We start with settings within the grid world domain, where modifications are limited to blocking cells and agent behavior is assumed to be optimal.
+This is the standard setting for the majority of goal recognition design studies, as highlighted in Table 1 of the survey by Keren et al.
+(2021).
 The objective of this set of simulations is to enable comparisons with state-of-the-art methods.
@@
 Environments where the goals are not reachable from the starting grid are filtered out.
-We randomly generate 200 environments and compare the average performance of our approach with that of baseline methods.
+We randomly generate at least 500 environments and compare the average performance of our approach with that of baseline methods.
+More details of data generation and wcd predictive model training are provided in Appendix A.1 and C.
+During optimization for our approach, the budget cannot be directly specified, so we explored over a range of Lagrangian parameters on a log scale between 0 and 10 (detailed in Appendix subsection B.2) and record the resulting realized budgets.
+We use the realized budget in our experimental figures because there is a monotonic relationship: larger Lagrangian parameters correspond to stricter penalties, generally producing lower realized budgets.
+This allows for a clearer comparison between methods that set budgets directly and our approach, which controls budgets indirectly through Lagrangian parameters.
 (a) $6\times 6$ grids
@@
 Exhaustive search and Pruned-Reduce are not included in (b) because they take more than an hour to compute for a single environment.
-(c) Shared budget
-(d) Individual budget
+We begin our experiments with a $6\times 6$ grid world.
+In this simplest setting, our approach and all baselines achieve similar performance in reducing wcd, as shown in Figure 3(a).
+However, our method demonstrates a significant run-time advantage over exhaustive search.
+In particular, our approach requires only 0.2 seconds, compared to approximately 2 seconds for exhaustive search.
+We then scale up the environment to a $13\times 13$ grid, where both exhaustive search and the Pruned-Reduce method failed to complete within an hour for a single instance and were therefore excluded from the baselines.
+As shown in Figure 3(b), our approach outperforms the greedy baselines in reducing wcd, due to its budget-aware optimization procedure that systematically accounts for the long-term impact of each modification—making it less prone to shortsighted or myopic decisions.
+Additionally, our method is approximately three times faster than the greedy baselines for large budgets (with further runtime comparisons discussed later).
+(a) Shared budget
+(b) Individual budget
 Figure 4: wcd reduction in settings with two types of modifications.
-We only included greedy baselines as the state-of-the-art baselines, such as Exhaustive search and Pruned-Reduce, are not applicable in these settings.
-We start our experiments with a grid world of size $6\times 6$.
-In this simplest scenario, our approach and all baselines exhibit similar performance in wcd reduction, as depicted in Figure 3(a).
-For runtime comparison, our method demonstrates a significant speed advantage over exhaustive search, performing on average 10 times faster.
-With the maximum allowed budget of 19 modifications, our approach required only 0.2 seconds, whereas the exhaustive search took approximately 2 seconds.
-[^2] Detailed runtime comparisons are included in the appendix.
-In this simplest setting, the runtime of our approach roughly matches that of the greedy method but is 10 times faster than exhaustive search.
-In all other settings, our method is several orders of magnitude faster than both exhaustive search and greedy using true wcd.
-Although the greedy baseline becomes much faster when utilizing our predicted model for wcd, our approach remains over 3 times faster than the greedy baseline with predicted wcd for large budgets.
-We then expanded the grid size to $13\times 13$.
-In this more complex setting, both exhaustive search and the Pruned-Reduce method were unable to complete computations within an hour for an instance, leading to their exclusion from the baselines.
-As illustrated in Figure 3(b), our approach surpasses the greedy baselines in wcd reduction and is 3 times faster than these baselines for large budgets.
-Figure 5: Overcooked-AI Environment. Assuming optimal agent behavior.
-Figure 6: Grid world ($6\times 6$). Incorporating suboptimal agent behavior.
-5.1.2 Settings with flexible budget constraints
+We only included greedy as the state-of-the-art baselines - Exhaustive search and Pruned-Reduce are not applicable in these settings.
+Flexible budget constraints.
 In the literature, most works focus on a single type of environment modification (e.g., blocking a cell).
-Given the flexibility of our optimization framework, we extend our approach to also include ’unblocking’ blocked cells as a possible environment update.
+Given the flexibility of our optimization framework, we extend our approach to include ’unblocking’ blocked cells as a possible environment update.
 In our simulations, we examine two common cases.
-In the shared budget setting, the total number of blocking and unblocking actions is bounded by a given shared budget.
-In the individual budget setting, the number of blocking and unblocking actions is bounded by different budgets.
+In “shared budget”, the total number of blocking and unblocking actions is bounded by a given shared budget.
+In “individual budget”, we limit the number of blocking and unblocking actions separately.
 Specifically, we allow the number of blocking actions to be 5 times the number of unblocking actions (rounded to the nearest integer).
-Given that there are no established baselines for this setting in the literature, we compare our results against greedy baselines.
-The results, as shown in Figure 4, demonstrate the effectiveness of our approach in wcd reduction.
-Regarding runtime, our approach is several orders of magnitude faster than the greedy method with true wcd and at least 3 times faster than the greedy method with predicted wcd.
-5.1.3 Complex domain and suboptimal agent behavior
-We consider two additional extensions to the standard setting.
-In the first, we evaluate the performance of our approach in a more complex problem domain: Overcooked-AI.
+Given that there are no established baselines (Pruned-Reduce is limited to blocking only modifications, and Exhaustive Search does not scale) for this setting in the literature, we compare our results against greedy baselines.
+The results are shown in Figure 4.
+Overall, they demonstrate that our approach is effective at reducing wcd even when existing state-of-the-art methods are not applicable.
+In addition, our gradient-based optimization outperforms the greedy approach except for small-budget cases [^2] Note that by definition of greedy approaches, the greedy approach using the true wcd is optimal when the budget is 1, since only one modification can be made.
+As a result, greedy methods generally perform well when the budget is very small, but our approach starts to outperform greedy with larger budgets..
+This is because our method considers how to distribute the budget over time, rather than selecting the highest-impact modification at each step without considering future opportunities.
+Complex domain and suboptimal agent behavior.
+We consider two additional extensions, where existing approaches in the literature do not apply, to demonstrate the flexibility of our framework.
+In the first, we evaluate our approach in a more complex problem domain: Overcooked-AI.
 In the second, we return to the grid world but include scenarios with suboptimal agent behavior.
@@
 Therefore, we compare our results with the greedy baselines.
-The results for both extensions, as illustrated in Figures 6 and 6, demonstrate that our approach adapts well to both settings.
+(a) Overcooked-AI
+(b) suboptimal agent.
+Figure 5: Performance in other settings: (a) Overcooked-AI environment with optimal agent behavior, and (b) Grid world ($6\times 6$) with suboptimal agent behavior, demonstrating the method’s adaptability to non-optimal decision-making.
+The results for both extensions, as illustrated in Figures 5(a) and 5(b), demonstrate that our approach adapts well to both settings.
+Our approach demonstrates a more significant advantage over baselines in Overcooked-AI than in grid world settings.
+The increased complexity and larger modification space in Overcooked-AI cause greedy methods to converge to short-sighted solutions, underscoring the superior scalability of our method.
 Regarding runtime, our approach is again several orders of magnitude faster than the greedy method with true wcd and at least 3 times faster than the greedy method using predicted wcd.
+Running time comparisons.
+We have demonstrated that our approach achieves greater wcd reductions compared to all baselines.
+Another key benefit of our method is its significantly improved runtime efficiency relative to standard approaches.
+This efficiency comes from (1) using a machine learning oracle to predict wcd instead of evaluating the true wcd at each time step, and (2) applying gradient-based optimization rather than the (heuristic-enhanced) exhaustive search methods used in the literature.
+As shown in our earlier simulations, exhaustive search methods scale poorly.
+In the 13x13 environment, a single instance of GRD takes more than one hour, while our approach completes in about one second.
+Therefore, we focus on the improvements due to the ML oracle in the following comparisons.
+Table 1: Time (seconds) comparisons when using true wcd and predicted wcd. $\pm$ denotes standard error.
+Setting True wcd Predicted wcd
+greedy greedy our approach
+[…]
+[…]
+[…]
+[…]
+Table 1 shows the time required to make modifications under a fixed budget across different settings.
+The results indicate our methods are generally faster and scale better in complex environments.
+For instance, in the complex domain, computing the true wcd takes 531 seconds for GRD, whereas our approach takes less than 1 second.
+Additional discussion on the time improvements is provided in Appendix A.2.
+5.1.1 Ablation Study
+We conducted an ablation study to better understand the specific contributions of our proposed approach to goal recognition design (GRD), particularly its ability to reduce wcd.
+The study isolates two key components of our method: the accuracy of the predictive model used to estimate wcd and the effectiveness of the optimization procedure in modifying the environment to improve goal recognition.
+We evaluated several predictive models trained to estimate wcd and tested them in combination with different optimization strategies.
+These are compared against a baseline that applies a greedy search using the true wcd without relying on a predictive model.
+All experiments were conducted under three environment modification settings: the standard setting and two variants with flexible budget constraints (individual and shared budgets).
+Results for environments with grid size 13 are shown in Table 2, with additional results for grid size 6 provided in the appendix.
+Impact of predictive model choice.
+We trained several predictive models, including Convolutional Neural Networks (CNNs), linear models, and Transformers, using the Adam optimizer (Kingma & Ba, 2015) and selected the configurations through validation.
+CNNs consistently achieved the lowest prediction error, and crucially, enabled more effective environment modifications when combined with either greedy or gradient-based optimization.
+By contrast, linear models and Transformers performed worse in predicting wcd and led to less effective downstream performance.
+This highlights the importance of predictive accuracy, where the quality of the predicted wcd directly influences the success of environment design.
+Impact of optimization approach.
+We also analyzed the role of the optimization procedure given a fixed predictive model, using (i) greedy optimization using predicted wcd, and (ii) our proposed gradient-based optimization approach.
+Greedy optimization using predicted wcd can offer runtime benefits over using true wcd, but its effectiveness varies with prediction quality.
+In cases where the predictor was less accurate, the greedy method often failed to produce significant improvements in wcd.
+Our proposed gradient-based approach consistently improved upon the greedy method.
+In particular, when combined with the most accurate CNN predictor (100K samples at a learning rate of 0.001), our method achieved the lowest wcd across all three modification settings, significantly beating the greedy baseline using true wcd as well.
+The best-performing configuration, CNN (100K, 0.001) combined with our gradient-based optimization, achieved the greatest reduction in wcd and outperformed all baseline approaches.
+This supports our core claim: accurate prediction of wcd paired with an effective optimization strategy, as adopted in our approach, is critical to designing environments that enhance goal recognition.
+Table 2: Ablation results showing the impact of different predictive models and optimization methods on wcd reduction across three environment settings as described in subsection 5.1.
+Each row represents a predictive model trained with a specified dataset size, learning rate, and associated training/validation loss.
+Due to poor performance, only the best Linear and Transformer configurations are shown.
+For each model, we compare two optimization methods: Greedy using predicted wcd and our proposed approach, along with a baseline using greedy with true wcd.
+Bolded results indicate the best performance per setting.
+Results are for grid size 13; grid size 6 results are in Appendix B.3
+Predictive Model Loss (mse) Optimization Approach wcd Reduction
+Train Validation Standard Individual Shared
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 5.2 Real-World Human-Subject Experiments
 In our simulations, we demonstrate that our approach consistently outperforms baseline methods in terms of wcd reduction and offers greater efficiency in runtime.
-To assess the applicability of our approach in settings where humans are the decision-makers, we conducted two sets of experiments involving human subjects.
-In the first experiment, our goal is to collect human behavioral data within our decision-making environments.
+To assess the applicability of our approach when humans are the decision-makers, we conducted two sets of human-subject experiments.
+In the first, we aim to collect human behavioral data in our decision-making environments.
 Utilizing this data, we employ imitation learning to develop a model that accurately represents human behavior.
@@
 5.2.1 Experiment 1: Collection of Human Behavioral Data
-In our experiment, we recruited participants to play 15 navigation games within a $6\times 6$ grid world.
-In each game, participants were tasked with navigating from a start position to a designated goal.
+We recruited 200 Amazon Mechanical Turk workers, paying each \$1.00 for an average task time of 3.64 minutes, equivalent to an hourly rate of about \$16.
+Each participant is asked to play 15 navigation games in a $6\times 6$ grid world, navigating from a start position to a designated goal in each game.
 At each time step, they could choose to move in one of four directions: Up, Down, Right, or Left.
@@
 Our objective with this setup was to leverage the collected data to develop a data-driven model of human behavior.
-We recruited 200 workers from Amazon Mechanical Turk for this study, paying an average hourly rate of approximately \$16 per worker.
 Learning models of human behavior.
-The collected human data were divided into training, validation, and testing sets: the training set included data from 160 workers, with approximately 70,000 instances of user decisions, while the validation and testing sets each contained data from 20 workers, with around 8,800 instances of user decisions each.
-We employed a 4-layer Multilayer Perceptron (MLP) model for training the model.
-The input to the model is the current environment layout, and its output predicts the next human action.
-We fine-tuned the hyperparameters based on the performance on the validation dataset.
-We compared the performance of our learned model against a model that assumes optimal agent behavior, which is defined as taking the shortest path to the goal.
-The training, validation, and test accuracies of both models are presented in Table 1.
-These results clearly indicate that human behavior deviates significantly from optimality.
-This deviation underscores the importance of incorporating a realistic model of human behavior in goal recognition design, particularly when humans are the decision-makers.
-Table 1: Prediction accuracy of human behavior assuming optimal behavior versus data-driven model.
-Assuming Optimal Behavior Using Data-Driven Model
-Training accuracy 0.7266 0.9189
-Validation accuracy 0.6964 0.8136
-Testing accuracy 0.7131 0.8422
+The collected human data were split into training (160 workers,  70,000 user decisions), validation, and testing sets (20 workers,  8,800 decisions each).
+A 4-layer Multilayer Perceptron (MLP) was trained, with the environment layout as input to predict the next human action.
+We fine-tuned hyperparameters and chose the model architecture using validation.
+We compared the performance of our learned model against one that assumes optimal agent behavior i.e., taking the shortest path to the goal.
+Table 3 presents both models’ accuracies, showing that human behavior significantly deviates from optimality.
+This deviation underscores the importance of incorporating a realistic model of human behavior in GRD, particularly when humans are the decision-makers.
+Table 3: Prediction accuracy of human behavior assuming optimal behavior versus using a data-driven model.
+Model Train Val. Test
+Assuming optimal behavior 0.7266 0.6964 0.7131
+Data-driven model 0.9189 0.8136 0.8422
 5.2.2 Experiment 2: Evaluating Goal Recognition Design
-We next evaluate the performance of our approach that incorporates the data-driven model of human behavior from experiment 1.
-To do this, we randomly generate 30 initial environments using the same setup as in our simulations.
+We next evaluate our approach that incorporates the data-driven model of human behavior from Experiment 1.
+We randomly generate 30 initial environments using the same setup as in simulations.
 These environments are then updated according to four different methods, all operating within a modification budget of 20.
@@
     Proposed (data-driven): Our proposed approach using predicted wcd from the data-driven model.
-We recruited 200 workers from MTurk.
+We recruited 200 workers from Amazon Mechanical Turk.
 Each worker was randomly assigned to one of the four treatments above, with the distinction between treatments being the environments presented to them.
 Workers were randomly assigned a goal for each environment and tasked with navigating to reach it.
-We utilize the collected data to evaluate different goal recognition design approaches.
+We utilize the collected data to evaluate GRD design approaches.
 (a) Empirical overlapping actions (the lower the better).
 (b) Bayesian goal inference (the higher the better).
-Figure 7: Comparing different goal recognition design approaches in our human subject experiments.
-Our approach coupled with data-driven models is shown to generate environments that enable the most effective goal recognition with real human decision makers.
+Figure 6: Comparing different GRD approaches in our human subject experiments.
+Our approach coupled with data-driven models is shown to generate environments that enable the most effective goal recognition with real human decision-makers.
 Comparing empirical overlapping actions.
-To evaluate the effectiveness of each goal recognition design approach, we first measure the empirical overlapping actions towards each of the two goals.
+To evaluate the effectiveness of each GRD approach, we first measure the empirical overlapping actions toward each of the two goals.
 Specifically, each recruited worker is exposed to all 30 environments in their assigned treatment and is instructed to reach one of two randomly selected goals.
-For each treatment, we calculate the number of overlapping actions for every pair of workers assigned to different goals within each environment.
-Therefore, for each treatment, we obtain a distribution of the number of overlapping actions.
-This metric reflects the difficulty of inferring agent goals for the environment and serves as an empirical proxy for wcd.
-Figure 7(a) presents the percentiles of the number of overlapping actions for each treatment.
-In particular, in the lowest 25%, the environments by our approaches have no overlapping actions, indicating that the environment makes it easy to infer the goal of the agents from observations.
-Generally, our approach, when integrated with data-driven models, leads to environments with fewer number of overlapping actions and therefore facilitate the easier identification of the humans’ intended goal.
+For each treatment, we compute the number of overlapping actions for every pair of workers assigned to different goals within the same environment.
+This yields a distribution of overlapping actions for each treatment.
+This metric reflects the difficulty of inferring an agent’s goal and serves as an empirical proxy for wcd.
+Figure 6(a) presents the percentiles of the number of overlapping actions across treatments.
+Notably, in the lowest 25%, environments generated by our approach show no overlapping actions, indicating that goal inference is relatively easy in these settings.
+Overall, when combined with data-driven models, our approach produces environments with statistically significantly fewer overlapping actions (paired t-test, p-value < 0.005), thereby facilitating more accurate identification of human goals.
 Comparing the accuracy of goal inference.
-Instead of using wcd or proxy of wcd for evaluations, we next directly measure whether we can indeed infer the human goals based on partial observations of their actions, utilizing an off-the-shelf Bayesian inference algorithm.
-[^3] We assume an initial uniform prior distribution.
-As the agent takes actions, we update the posterior belief over goals based on the likelihood probability of the data-driven behavior model prediction.
-Specifically, for each worker, we reveal the first $k$ portion of the worker’s actions to the inference algorithm to infer the worker’s goal.
-This enables us to compute the average inference accuracy.
-The results in Figure 7(b) demonstrate that our approach leads to environments that are easier for goal recognition.
-6 Conclusion
-Effective human-AI collaboration hinges on the AI’s ability to infer the goals of humans.
-In this study, we work on the problem of goal recognition design, updating the decision-making environments to make it easier for the AI agents to perform goal recognition.
-By addressing the key limitations in the existing literature, notably the computational demand and the assumption of optimal agent behavior, we have developed a data-driven optimization framework that is both efficient and scalable.
-Through simulations and human-subject experiments, we show that our approach outperforms state-of-the-art approaches in standard settings, applies to settings that existing approaches cannot address, and leads to environments easier for goal recognition with real-world human decision makers.
-Societal impact.
-Our approach is capable of integrating general models of various agent behavior, making it more applicable to real-world scenarios involving human agents.
-However, with more accurate human models, it also opens up the concerns of privacy and potential mis-use of the techniques for unethical surveillance.
-While our work does not directly facilitate such misuse, the broader field of goal recognition research could potentially be exploited.
+Instead of evaluating the effectiveness of our approaches based on wcd or its proxy, we next directly assess whether human goals can be accurately inferred from partial observations of their actions.
+To do this, we use an off-the-shelf Bayesian inference algorithm (Murphy, 2012)  [^3] We assume a uniform prior distribution over goals.
+As the agent takes action, we update the posterior belief over goals through Bayesian updates based on the likelihood probability of the data-driven behavior model prediction..
+As the agent acts, we update the posterior belief using the likelihood derived from our data-driven behavior model.
+Specifically, for each worker, we reveal the first $k$ portion of their actions to the inference algorithm, which then attempts to infer the worker’s goal.
+This allows us to compute the average inference accuracy.
+The results, shown in Figure 6(b), demonstrate that our approach produces environments that make goal recognition easier.
+6 Conclusion and Future Work
+Effective human-AI collaboration hinges on the AI system’s ability to infer human goals.
+In this work, we introduce a data-driven framework for goal recognition design (GRD) that optimizes environments to enhance goal disambiguation.
+Our method addresses two major limitations of prior approaches: computational inefficiency and the unrealistic assumption of fully optimal agent behavior.
+At the core of our framework is a learned oracle that predicts goal recognition difficulty, worst-case distinctiveness (wcd).
+By integrating this oracle into a gradient-based optimization process, we avoid expensive exact evaluations and unlock a scalable, flexible design procedure.
+This enables three core advantages: (1) improved scalability, (2) support for general, potentially suboptimal agent models, and (3) extensibility to diverse constraints and objectives.
+Our empirical evaluation highlights the effectiveness and versatility of the approach.
+In both classical grid worlds and complex Overcooked-AI environments, our method consistently outperforms baseline techniques in reducing wcd while significantly lowering computational cost.
+Crucially, it remains effective under suboptimal agent behavior—demonstrating robustness where many existing methods break down.
+Human-subject experiments further validate the practical utility of our framework, showing that it generalizes beyond simulation and maintains its advantage even when agent models are learned from real human data.
+To our knowledge, this is the first work to evaluate GRD methods in a human-subject setting.
+Through ablation studies, we demonstrate that the predictive oracle plays a critical role in guiding effective environment optimization, with more accurate or expressive models leading to significantly better outcomes.
 Limitations and future work.
-While our work has extended the literature in goal recognition design to address suboptimal agent behavior, more flexible budget constraints, and more complicated environments, there are still several limitations.
-In particular, we have assumed static and full observability of the environment.
-Extending our framework to deal with more complex environments, particularly those that are dynamic and partially observable would be an interesting and important future work.
-Moreover, we have assumed that user behavior remains static, so we can learn a human model from their historical behavior.
-Finally, given our framework is adaptable to different forms of agent behavior, understanding how different patterns of agent suboptimality impacts the difficulty of goal recognition could lead to useful insights for enabling better human-AI collaboration.
-Acknowledgements
-This work is supported in part by J.P. Morgan Faculty Research Award and a Global Incubator Seed Grant from McDonnell International Scholars Academy.
+While our framework advances goal recognition design, several limitations remain.
+First, our approach requires a large amount of data to train accurate predictive oracles.
+Although simulations can be used to generate training data, data from novel and richer environments is often scarce and presents challenges for simulation.
+Future work could explore several methods to address this limitation, including leveraging transfer learning (Pan & Yang, 2009) by pretraining on simulated environments and fine-tuning on limited data from new environments, as well as using generative models to produce synthetic yet behaviorally plausible training data (Goodfellow et al., 2014; Xu et al., 2019).
+Active learning techniques that reduce the amount of data required during collection also offer a promising direction.
+While these strategies require further validation, we believe they represent potentially viable and interesting avenues for applying GRD in data-scarce settings.
+Second, we assume fully observable environments and stationary human behavior in this work.
+To address these limitations, future work could train the predictive model on inputs derived from belief state representations, incorporating uncertainty in agent observations (Kaelbling et al., 1998) into the training process.
+In addition, the predictive model could be periodically fine-tuned using newly accumulated behavioral data, allowing it to adapt to changes in human decision-making over time.
+Moreover, our investigation focuses on single-agent, discrete environments.
+To extend our framework to multi-agent settings, future work would need to model agent behavior in the presence of other agents, accounting for their interactions.
+To support continuous environments, the Lagrangian relaxation can be adapted to penalize changes using continuous distance functions such as norms, enabling gradient-based optimization without altering the underlying pipeline.
+Overall, the modularity of our framework enables future work to address these limitations in a targeted and flexible manner.
+Broader impact. Finally, while our work aims to advance human–AI collaboration through goal recognition design, it is important to acknowledge the societal risks that may arise from its misuse.
+First, designing environments that make human goals and intentions easier to infer could be misused.
+Systems capable of anticipating intentions might be deployed not to assist users but to manipulate or steer their behavior.
+Because system designers typically hold an advantageous position over decision makers within these environments, such capabilities can further amplify their influence over individuals.
+Second, since GRD relies on predictive models trained on human behavioral data, it raises privacy concerns.
+Even when raw data are not directly exposed, adversaries could analyze model outputs or behaviors to infer sensitive personal attributes, creating opportunities for profiling or surveillance.
+Third, by making goal inference more accurate and efficient, GRD could foster excessive reliance on AI predictions.
+Such overreliance risks undermining human autonomy and weakening critical judgment, especially if users begin to accept AI inferences uncritically.
+Taken together, these risks highlight how a technique intended to enhance collaboration could, under certain deployments, compromise user welfare and agency.
+In light of these concerns, we discuss possible avenues for risk mitigation.
+On the technical side, privacy-preserving methods such as differential privacy could be integrated into the training of behavioral models, limiting the extent to which sensitive information is encoded and thereby reducing opportunities for exploitation.
+Moreover, incorporating uncertainty-aware methods into GRD could ensure that systems explicitly communicate the confidence of their inferences.
+Such mechanisms would help mitigate both privacy risks and overreliance by reminding users of the fallibility of AI systems.
+On the policy side, once the capabilities and risks of GRD are more fully understood, it will be important to weigh the potential benefits against the possible harms.
+This evaluation should guide the development of regulatory frameworks and ethical guidelines—for example, requiring that GRD-driven modifications not significantly reduce user autonomy relative to standard environments, and mandating transparency about the objectives and data underlying deployment.
+Establishing such safeguards will be essential to ensure that GRD is developed and applied responsibly, and that its benefits are realized without compromising broader societal values.
+Acknowledgments
+This work is supported in part by a J.P. Morgan Faculty Research Award and a Global Incubator Seed Grant from the McDonnell International Scholars Academy.
 Appendix A Experiment Details and Additional Results
-In this section, we provide details of the experiments that have not been included in the main paper due to space constraints. We also include and discuss additional experiment results.
+In this section, we provide details of the experiments not included in the main paper due to space constraints. We also include and discuss additional experimental results.
 A.1 Implementation Details
-In this subsection, we discuss the details of our implementation for training the predictive model for wcd.
-To train a CNN-based prediction model, we generated an extensive training dataset by solving wcd for a large number of randomly selected environments, given a specific agent model.
-For creating a predictive model for environments with a grid size of $N$ by $N$, we utilized a custom Convolutional Neural Network (CNN) architecture.
-This architecture is tailored to process input data, $w_{i}$, consisting of $k$ $N\times N$ channels, where $k$ represents the number of potential objects in the environment.
-The architecture comprises two distinct blocks: the initial block contains three successive convolutional layers, each followed by batch normalization and a rectified linear unit (ReLU) activation function.
-The subsequent block introduces pooling operations and additional convolutional layers for more sophisticated feature extraction.
-After the convolutional blocks, the architecture includes fully connected layers, leading to a single output unit.
-The design integrates Leaky ReLU, ReLU, and dropout layers to add non-linearity and provide regularization.
-This flexible architecture can adapt to tasks with multi-channel input data, allowing for modifications based on the specific needs of the task and dataset characteristics.
-Batch normalization and dropout are incorporated to promote training stability and prevent overfitting.
-We employed the Adam optimizer and Mean Squared Error (MSE) loss for training the models, specifically focusing on settings involving optimal human agents.
-To create the dataset to train the model, we randomly generated $150$K grid designs and computed the corresponding wcd.
-This dataset was then split into the training dataset ($80\%$) and validation dataset ($20\%$).
-For our experiments reported in Section 5 with $13\times 13$ grid size and optimal agent behavior, we have achieved an MSE value around 0.18 for both training and validation errors.
-The results are shown in Figure 8.
-We have also examined a variety of setups, and generally we can reach small errors for the predictive model that can enable the optimization.
-Figure 8: Training loss and validation errors for the predictive model for wcd with $13\times 13$ grid world and optimal human behavior.
-A.2 Results for Run Time Comparisons
+This subsection describes the implementation of the predictive model.
+We trained the predictive models on a simulated dataset generated by solving wcd for randomly sampled environments with a specific agent behavior model.
+Several architectures were tested, including CNNs, Linear Models, Kernel Ridge Regression, and Transformers, with CNNs performing the best in predicting wcd.
+The models processes inputs of the shape $k\times N\times N$ inputs, where $k$ is the number of potential objects in an $N\times N$ grid.
+Figure 7: Training loss and validation errors for the CNN predictive model for wcd with $13\times 13$ grid world and optimal human behavior.
+The CNN architecture used across the paper employs a ResNet18 backbone with an adapted $k$-channel input layer, followed by three fully connected layers.
+Outputs are reduced to 32 and 16 dimensions with Leaky ReLU activations and dropout, culminating in a single non-negative scalar via ReLU activation.
+For the experiments in Section 5, we trained a CNN-based model on 100K data points, split into 80% training and 20% validation.
+The model achieved an MSE below 0.18 for both sets, with the validation set used for hyperparameter tuning.
+Figure 7 shows the training loss for the wcd predictive model for the grid world domain ($13\times 13$) assuming optimal behavior.
+We used Adam optimizer and MSE loss and tested learning rates of 0.1, 0.01, 0.001, and 0.0001.
+A learning rate of 0.001 consistently produced the lowest validation error, enabling reliable wcd reduction across different setups.
+A.2 More Results for Run Time Comparisons
 Due to space constraints and also that the results align with one would expect, we do not include the run time details for different approaches in the main text.
@@
 Figure 8: Run time taken by each method in the different experimental conditions.
-In the standard setup with a small grid size, all methods except exhaustive complete within less than half of a second but only our approach scales with a larger grid size and other more complex configurations.
-All our experiments are conducted on a cluster of 40 CPU cores (Intel Xeon Gold 6148 CPU @@@@2.40GHz), 1 GPU (NVIDIA Tesla V100 SXM2 32GB), and a maximal memory of 80GB.
+In the standard setup with a small grid size, all methods except exhaustive complete within less than half of a second but our approach scales better with more complex configurations.
 Appendix B Additional Experiment Results
 In this section, we report the additional experiment results that are not included in the main text due to space constraints.
-In Section 5.1.2 of the main paper, we provided details of our performance with a large grid size.
+In Section 4 of the main paper, we provided details of our performance with a large grid size.
 In a smaller grid world, our approach achieves comparable performance to the baselines but it significantly outperforms them in larger grid sizes as shown in Figure 9.
@@
 The budget in this domain is quantified as the total Manhattan distance moved by the objects between the original and modified environments.
+B.2 Langrange Parameter Space
+In the main paper, we described how we relax the original constrained optimization problem by introducing a Lagrangian penalty parameter $\lambda$, which indirectly controls the budget through penalizing constraint violations.
+Since the budget cannot be specified directly, we instead vary $\lambda$ across a predefined range to explore different trade-offs between optimizing the objective and satisfying the budget constraint.
+Specifically, we sweep over a log-scaled set of values:, $\lambda\in\{0,0.001,0.002,0.005,0.007,0.01,0.02,0.05,0.07,0.1,0.2,0.5,0.7,1.0,2,5,7\}$.
+This range covers several orders of magnitude, allowing the optimization to balance constraint enforcement from very loose (small $\lambda$) to very strict (large $\lambda$) penalties.
+For all experiments, including those with a single budget constraint, we used this same set of $\lambda$ values.
+In scenarios involving two constraints, such as Blocking and Unblocking constraints, we applied this range independently to each constraint, effectively exploring the joint penalty space by combining these $\lambda$ values for each constraint.
+B.3 Ablation Analysis
+In the main paper, we provided the ablation analysis for the optimal behavior model with a grid size of 13.
+Table 4 provides the same analysis for grid size 6.
+The results show mixed performance in this smaller environment, and our approach sometimes underperforms baselines.
+This reflects that our method’s advantages emerge primarily in larger, more complex environments where traditional approaches become computationally intractable.
+[…]
+Predictive Model Loss (mse) Optimization Approach wcd Reduction
+Train Validation Standard Individual Shared
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Appendix C Details of Simulations
-To generate our training and evaluation datasets, we randomly generated environments and kept environments that are valid, e.g., the goals are reachable, and that the objects don’t overlap.
-Below we provide more details for environment generation for specified grid sizes for both overcooked and grid-world domains.
+To generate our training and evaluation datasets, we randomly generated environments and kept valid environments, e.g., those in which the goals are reachable and the objects don’t overlap.
+Below, we provide more details about environment generation for specified grid sizes for both overcooked and grid-world domains.
 C.1 Grid world
@@
 The starting position is randomly placed in the first column, and the goal positions are randomly placed in the last two columns.
-The number of blocked positions in each grid is randomly selected from a range of 00 to 2×2$\times$2 × the grid width.
+The number of blocked positions in each grid is randomly selected from a range of 0 to $2\times$ the grid width.
 We discard any assignments that make the goals unreachable.
 For experiments with suboptimal behavior, we also randomly assigned small subgoal rewards to unblocked cells that the agent would collect on its way to the goal state.
-The two goal states assigned a large reward that is 10 times the largest subgoal reward.
+The two goal states are assigned a large reward that is 10 times the largest subgoal reward.
 C.2 Overcooked-AI
@@
 All objects must be reachable from the agent’s randomly assigned starting position, with the number and positions of blocked cells assigned randomly.
-The agent is randomly assigned any two possible goals: three tomato soups, three onion soups, or a mixed soup.
+The agent is randomly assigned one of two goals for each experiment, where both goals are randomly selected from the set of three possible recipes: three-tomato soup, three-onion soup, or mixed soup.
 Each goal has the same randomly assigned reward value.
@@
 Other 2
+Appendix E Computation Details
+All experiments were run on a computing cluster equipped with 40 CPU cores (Intel Xeon Gold 6148 @ 2.40GHz), a single NVIDIA Tesla V100 SXM2 GPU (32GB), and up to 80GB of memory.
+To ensure reproducibility, we worked within a clean and consistent software environment built around Python 3.10 and widely used scientific libraries.
+PyTorch was our main deep learning framework, with NumPy and pandas handling numerical computation and data processing.
+During model training and the discrete gradient descent optimization, we relied on PyTorch’s autograd system to compute gradients automatically and efficiently.